### v1

# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

Abstract
In real-world scenarios, time series forecasting often demands timeliness, making research on model backbones a perennially hot topic. To meet these performance demands, we propose a novel backbone from the perspective of information fusion. Introducing the Basic Probability Assignment (BPA) Module and the Time Evidence Fusion Network (TEFN), based on evidence theory, allows us to achieve superior performance. On the other hand, the perspective of multi-source information fusion effectively improves the accuracy of forecasting. Due to the fact that BPA is generated by fuzzy theory, TEFN also has considerable interpretability. In real data experiments, the TEFN partially achieved state-of-the-art, with low errors comparable to PatchTST, and operating efficiency surpass performance models such as Dlinear. Meanwhile, TEFN has high robustness and small error fluctuations in the random hyperparameter selection. TEFN is not a model that achieves the ultimate in single aspect, but a model that balances performance, accuracy, stability, and interpretability. Code is available at this repository: https://github.com/ztxtech/Time-Evidence-Fusion-Network.
1 Introduction
The changes of things are accompanied by the passage of time, and time series appear in more and more application scenarios, such as health diagnosis, financial analysis, weather forecasting, and so on. Optimizing structure of backbone is an effective way to improve the performance of time series forecasting, where linear-based models such as Dlinear often outperform transformer-based models such as Crossformer [18, 20]. Therefore, we hope to propose a model that balances performance, prediction accuracy, stability, and interpretability to provide support for highly reliable scenarios with time requirements.
Evidence theory is a model of uncertain reasoning commonly used for decision-making in the fusion of multiple information [14, 6]. Meanwhile, evidence theory is often appeared in reasoning framework based on symbolic logic, with high interpretability. Denoeux has implemented evidence neural networks for classification tasks [8]. Therefore, we propose the Time Evidence Fusion Network (TEFN) for regression tasks. The strategy adopted by TEFN is to divide the time series into two perspectives: time and channels, treating each channel/time as an independent information source, and using an information fusion framework for prediction in Figure 1. Therefore, TEFN is a simple and naive model, without any technical modifications, and is a new backbone.
Figure 1: The channel dimension and time dimension of multivariate time series can be considered as different sources of information. Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
In TEFN, we propose a novel module Basic Probability Assignment (BPA) Module. Based on the evidence theory of decision-making process, BPA maps different sources of information to distributions related to the target outcome. The target distribution is achieved by the fuzzy membership of different fuzzy sets [7]. In image processing, convolution condenses the required features of the target [4]. Unlike convolution, BPA is an expansion process in Figure 2. On the contrary, BPA considers all possibilities in the event space composed of subsets of the sample space, expanding on different events, which is more effective for simple structures like time series and can fully explore the different hidden information inside. For different distributions generated by multiple information sources, we consider sources as samples and use the expected forecasting values as the final result.
Figure 2: Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability. At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects. TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20] or Patch [12]. For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model. The contributions of this article are as follows:
  • This article proposes a new evidential backbone for extracting potential distributions of simple data structures.
  • A neural network TEFN based on BPA for time series forecasting has been proposed, which leads previous research in accuracy, efficiency, stability, and interpretability.
  • This article explicitly introduces relevant concepts from evidence theory, such as mass function, into the structure of neural networks, providing a new implementation approach for uncertainty reasoning.
2 Related Work
2.1 Dempster-Shafer Theory
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [14, 6]. It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. When combining two mass functions, denoted as $m_{1}$ and $m_{2}$, the resulting fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation 1.
$m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
2.2 Evidence Decision Making
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [7]. BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. Finally, decision making transpires as the progression from the attained distribution towards the designated target.
3 Time Evidence Fusion Network
The structure of TEFN is naive, and well performance can be achieved through a structure with only $2$ BPA module similar to only $2$ convolution kernels. We believe that the new architecture of TEFN is more like a pure backbone model. The overall structure of TEFN is in Figure 3.
Figure 3: The overall structure of TEFN
3.1 Time Normalization and De-normalization
In order to accelerate converging to local optima more quickly, we referred to the normalization methods of Stationary in TEFN [11]. The normalization process calculates the mean $\mu$ in Equation 2 and variance difference $\sigma^{2}$ in Equation 3 for each piece of the time series $x$. The normalization and de-normalization processes are a pair of symmetric operations in Equation 4 and Equation 5, where $Net(\cdot)$ is the neural network operation performed on the normalized time series $x_{Norm}$.
$\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$ (2)
$\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$ (3)
$x_{Norm}=\frac{x-\mu}{\sigma}$ (4)
$x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (5)
3.2 Time Dimension Projection
In TEFN, a linear layer is used to project the length $L_{in}$ of the time dimension to length $L_{in}+L_{pred}$ where $L_{pred}$ is prediction length of current target.
3.3 Basic Probability Assignment Module
Figure 4: The operation process of a BPA module for a time series $T$ of length $L=5$. The sample space size in the BPA module is set as $|S|=2$, and the event space size is set as $|2^{S}|=2^{|S|}=4$. The operation of the BPA module is similar to that of the convolution module, but the BPA module diffuses different masses instead of summing them in the convolution.
According to Evidence Theory, each dimension of a time series can be reduced to a mass function in the same set of event spaces. The event space of evidence theory usually chooses power sets $2^{S}$ of sample space $S$. If it is a classification problem, the sample space is all the labels. For regression problems, the sample space needs to be treated as anonymous samples, where the samples are mutually exclusive. Similarly, whether it is a classification or regression problem, each dimension of the data will generate a mass distribution in the event space. Fuzzy numbers commonly used in the generation of mass distribution of evidence.
Here we use a parameterized fuzzy membership $\mu$ in Equation 6, where $D$ (time dimension $T$ and channel dimension $C$ are available) is the dimension that generates the mass function, $j$ is corresponding to dimension $D$ and $i$ is corresponding to the other one. Parameter $w$ and $b$ represent the slope and intercept of the membership function, and $k$ is corresponding to event dimension $F$.
$m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (6)
3.4 Expectation Fusion
For multivariate time series in shape $|T|*|C|$, there are two dimensions: time dimension $T$ and channel dimension $C$. So, the different mass distributions $m_{C},m_{T}$ can be generated simultaneously. And the fusion method here uses expectations in Equation 7 where $y_{j,k}$ is parameter and $m^{\prime}_{i,j,k}$ is corresponding mass function. Summing part $y_{j,k}*m^{\prime}_{i,j,k}$ can be regarded as a linear transform of $m_{i,j,k}$, and expectation is calculated by summing $m_{i,j,k}$ in Equation 7. There are two reasons why DSR are not used:
  1. 1.
    The mathematical operation corresponding to DSR is the point-to-point multiplication of mass distributions within the same dimension, which has a high computational complexity. At the same time, if the dimension is too high, it will lead to the disappearance of accuracy within a finite precision.
  2. 2.
    DSR is sensitive to one extreme distribution, such as a distribution with only one element being $1$, which can result in the final fusion result of the entire mass being directly equal to that distribution.
$\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$ (7)
At the end, results of time dimension $T$ and channel dimension $C$ will aggregate in Equation 8 where $\hat{y}_{norm}$ is normalized predicted results. The final prediction results are restored through the de-normalization layer.
$\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (8)
4 Experiments
Dataset
We will use TEFN for long-term prediction tasks. Table 1 shows the basic settings used in the experiment, including the datasets. In terms of datasets, we have selected $5$ common time series datasets, including Electricity [1], ETT (4 subsets) [10], Exchange [17], Traffic [2], Weather [3]. More dataset features are in the Appendix Table 5.
Table 1: Datasets and Basic Settings
Tasks
Datasets
Metrics
Series Length
Long-term
ETT (4 subsets), Electricity,
MSE, MAE 96$\sim$720
Forecasting
Traffic, Weather, Exchange
Baseline
The comparative experimental models include PatchTST (ICLR 2023) [12], Crossformer (ICLR 2023) [20], TiDE [5], TimeNet (ICLR 2023) [16], ETSformer [15], LightTS [19], Dlinear (AAAI 2023) [18], FEDformer (ICML 2022) [21], Stationary (NeurIPS 2022) [11] and Autoformer (NeurIPS 2021) [17] . The configurations of the comparative experiments are from the Time Series Library (https://github.com/thuml/Time-Series-Library). More implementation details are in the Appendix Section A.
4.1 Forecasting Result
TEFN partially achieved SOTA performance in long-term forecasting target in Table 2. It is worth noting that in the case of a simple structure with tiny parameters in Section 4.4, achieved approximate or even better performance than previous SOTA models with larger parameters, such as PatchTST [12]. Full forecasting results are in the Appendix Table 7.
Table 2: Prediction error. All results are the average of four predicted lengths $\{96,192,336,720\}$.
Models
TEFN
PatchTST
Crossformer
TiDE
TimesNet
ETSformer
LightTS
Dlinear
FEDformer
Stationary
Autoformer
[12]
[20]
[5]
[16]
[15]
[19]
[18]
[21]
[11]
[17]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.403 0.398 0.387 0.400
0.513
0.496
0.419
0.419
0.400 0.406
0.429
0.425
0.435
0.437
0.403
0.407
0.448
0.452
0.481
0.456
0.588
0.517
ETTm2
0.286 0.327 0.281 0.326
0.757
0.610
0.358
0.404
0.291 0.333
0.293
0.342
0.409
0.436
0.350
0.401
0.305
0.349
0.306
0.347
0.327
0.371
ETTh1
0.441 0.429
0.469
0.454
0.529
0.522
0.541
0.507
0.458
0.450
0.542
0.510
0.491
0.479
0.456 0.452 0.440
0.460
0.570
0.537
0.496
0.487
ETTh2
0.380 0.402 0.387 0.407
0.942
0.684
0.611
0.550
0.414 0.427
0.439
0.452
0.602
0.543
0.559
0.515
0.437
0.449
0.526
0.516
0.450
0.459
Electricity
0.215
0.292 0.205 0.290
0.244
0.334
0.251
0.344
0.192 0.295
0.208
0.323
0.229
0.329
0.212
0.300
0.214
0.327
0.193
0.296
0.227
0.338
Exchange
0.362
0.402
0.259 0.281 0.259 0.315 0.271 0.320
0.416
0.443
0.410
0.427
0.385
0.447
0.354
0.414
0.519
0.500
0.461
0.454
0.613
0.539
Traffic
0.623
0.372 0.367
0.404
0.940
0.707
0.370
0.413
0.620
0.336
0.621
0.396
0.622
0.392
0.625
0.383
0.610
0.376
0.624
0.340
0.628
0.379
Weather
0.260 0.283
0.481
0.304
0.550
0.304
0.760
0.473
0.259 0.287
0.271
0.334
0.261
0.312
0.265
0.317
0.309
0.360
0.288
0.314
0.338
0.382
$1^{\text{st}}$ Count
8 20 16 13
1
0
1
0
7 5
0
1
0
0
1
0
5
0
1
1
0
0
$2^{\text{nd}}$ Count 13 7 7 15
1
3
4
3
6 5
2
1
0
0
2
2
0
0
5
4
0
0
$3^{\text{rd}}$ Count
5 5
4
3
3
3
2
2
11 20
4
0
3
0
6 4
2
1
0
1
0
1
4.2 Hyperparameter Sensitivity
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 5. It is worth noting that the order of magnitude of the MSE in Figure 5 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. Full hyperparameter sensitivity experiment results are in the Appendix Table 8.
(a) $lr$ $|S|$MSE
(b) Violin diagram of $|S|$ and MSE
Figure 5: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table 3. The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation. Random parameter selection within the traversal range of TEFN yields stable prediction results.
Table 3: The variance of model error for traversal learning rate and sample space overhaul
Metric Dataset
($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
192 0.0001 0.0000 0.0003 0.0004 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0002 0.0004 0.0000 0.0023 0.0000 0.0001
720 0.0000 0.0000 0.0004 0.0001 0.0000 0.0038 0.0000 0.0001
MAE 96 0.0000 0.0000 0.0001 0.0001 0.0001 0.0000 0.0000 0.0002
192 0.0001 0.0000 0.0001 0.0001 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0001 0.0001 0.0000 0.0009 0.0000 0.0001
720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
4.3 Ablation Study
TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$. In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table 4. Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error. Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (error of Electricty only varies $0.001$). It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
Table 4: The results of the ablation experiment
Metric Models
Dataset
Electricity
ETTh1
ETTh2
ETTm1
ETTm2
Exchange
Traffic
Weather
MSE C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
MAE C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
4.4 Efficiency Comparison
Figure 6: Model efficiency comparison. Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
TEFN is a compact model compared to the parameters of the Transformer. In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison. In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample. Different models occupy varying amounts of GPU and memory. Therefore, we used the size of the saved binary model files as the model size. We visualized this with a bubble chart in Figure 6 . It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters. Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency. TEFN maintains almost Dlinear [18] performance and almost PatchTST [12] accuracy, making it a performance error balanced model.
4.5 Interpretability Analysis
TEFN will allocate mass based on the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, and mass indicates the confidence reflecting support of different possible events. This support allocation is interpretable because BPA allocation utilizes fuzzy membership $\mu(x)=wx+b$ in Figure 7. Essentially, different fuzzy sets are constructed through TEFN for inference.
(a) Channel BPA
(b) Time BPA
Figure 7: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time. The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
4.6 Limitation and Future Prospects
TEFN is a naive and unmodified model. We lack skills in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method. Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well. In addition, for the fusion of mass distribution in the face of the failure of the classic Dempster Shafer rule for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern. In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels. In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution. Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic. Effective parameter initialization may exist, for example, initializing corresponding parameters in a non neural network way may lead to faster convergence.
5 Conclusion
This paper treats different channels and different time points of time series as multiple sources of information. By utilizing evidence theory to construct different basic probability assignments, TEFN discards irrelevant time points and channels. Experimentally, TEFN is a parameter-insensitive and stable prediction model with a small parameter count and fast training speed. In future work, we will explore how to optimize basic probability assignments and information fusion in neural networks.
Acknowledgments
The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078).
Appendix A Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores , 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch [13]. For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The loss function selected during the experiment was MSE, and the optimizer selected Adam [9]. We used Mean Absolute Error (MAE) and Mean Square Error (MSE) in Equation 9 and Equation 10 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
$MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (9)
$MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (10)
Appendix B Supplementary Experimental Results
B.1 Dataset Feature
In the experimental section, we mainly referred to the experimental dataset and experimental settings of TimesNet. The experimental settings of TimesNet are comprehensive and reasonable [16]. Table 5 shows the quantitative features of the dataset used in the experiment. Dataset size corresponds to the training set, validation set, and test set. Each dataset has four forecasting tasks of different lengths $L\in\{96,192,336,720\}$.
Table 5: The quantitative features of the dataset
Dataset Dimension Dataset Size
Electricity 321 (18317, 2633, 5261)
ETTm1, ETTm2 7 (34465, 11521, 11521)
ETTh1, ETTh2 7 (8545, 2881, 2881)
Traffic 862 (12185, 1757, 3509)
Weather 21 (36792, 5271, 10540)
Exchange 8 (5120, 665, 1422)
B.2 Forecasting Visualization
We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models. We take the task of Electricity-96 [1] as an example and selected typical models PatchTST [12], Crossformer [20], TimesNet [16], Dlinear [18] and Stationary [11] for visualization in Figure 8. From the visualization results, TEFN fits the details and trends of time series better.
(a) TEFN (Ours)
(b) PatchTST
(c) Crossformer
(d) TimesNet
(e) Dlinear
(f) Stationary
Figure 8: Visualization of Electricity-96 tasks
B.3 Full Comparative Results
In the comparative experiment, we conducted a thorough comparison of 5 datasets and 10 comparative models in terms of forecasting length $l\in\{96,192,336,720\}$ in Table 7. It can be seen that TEFN has positive predictive performance on the experiment dataset, achieving SOTA in multiple forecasting tasks across multiple datasets.
Table 6: Full results of the ablation experiment
Metric Target Model Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
C 0.196$\downarrow$ 0.386$\uparrow$ 0.288 0.339$\downarrow$ 0.185$\uparrow$ 0.083$\uparrow$ 0.644$\downarrow$ 0.185$\uparrow$
T 0.197 0.386$\uparrow$ 0.291$\uparrow$ 0.356$\uparrow$ 0.182$\uparrow$ 0.083$\uparrow$ 0.645 0.195$\uparrow$
192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
C 0.196$\downarrow$ 0.435$\uparrow$ 0.375 0.382$\uparrow$ 0.246 0.179$\uparrow$ 0.597$\downarrow$ 0.230$\uparrow$
T 0.197 0.435$\uparrow$ 0.375 0.390$\uparrow$ 0.247$\uparrow$ 0.175$\downarrow$ 0.598 0.240$\uparrow$
336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
C 0.210$\downarrow$ 0.479$\uparrow$ 0.423 0.411$\downarrow$ 0.307 0.336$\uparrow$ 0.604$\downarrow$ 0.280$\uparrow$
T 0.212 0.479$\uparrow$ 0.414$\downarrow$ 0.423$\uparrow$ 0.307 0.324$\downarrow$ 0.606$\uparrow$ 0.291$\uparrow$
720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
C 0.253 0.478$\uparrow$ 0.457$\uparrow$ 0.472$\downarrow$ 0.407 0.947$\uparrow$ 0.642$\downarrow$ 0.354$\uparrow$
T 0.254$\uparrow$ 0.479$\uparrow$ 0.442$\uparrow$ 0.487$\uparrow$ 0.408$\uparrow$ 0.871$\uparrow$ 0.643$\downarrow$ 0.364$\uparrow$
Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
C 0.273$\downarrow$ 0.394$\uparrow$ 0.339$\uparrow$ 0.366$\downarrow$ 0.269$\uparrow$ 0.199$\uparrow$ 0.384 0.228$\uparrow$
T 0.274 0.393$\uparrow$ 0.339$\uparrow$ 0.377$\uparrow$ 0.265$\uparrow$ 0.199$\uparrow$ 0.383$\downarrow$ 0.234$\uparrow$
192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
C 0.275$\downarrow$ 0.421$\uparrow$ 0.393$\uparrow$ 0.385$\uparrow$ 0.305$\uparrow$ 0.300$\uparrow$ 0.361$\uparrow$ 0.264$\uparrow$
T 0.276 0.421$\uparrow$ 0.390$\downarrow$ 0.392$\uparrow$ 0.305$\uparrow$ 0.296$\downarrow$ 0.361$\uparrow$ 0.270$\uparrow$
336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
C 0.291$\downarrow$ 0.443$\uparrow$ 0.434$\downarrow$ 0.403$\downarrow$ 0.343 0.419$\uparrow$ 0.363 0.298
T 0.292 0.445$\uparrow$ 0.425$\downarrow$ 0.413$\uparrow$ 0.342$\downarrow$ 0.410$\downarrow$ 0.362$\downarrow$ 0.306$\uparrow$
720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
C 0.325 0.463$\downarrow$ 0.462$\uparrow$ 0.436$\downarrow$ 0.399 0.739$\uparrow$ 0.382 0.347$\uparrow$
T 0.324$\downarrow$ 0.469$\uparrow$ 0.454$\uparrow$ 0.449$\uparrow$ 0.398$\downarrow$ 0.702$\uparrow$ 0.382 0.354$\uparrow$
Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
Table 7: Full results of comparative results
Models
TEFN
PatchTST
Crossformer
TiDE
TimesNet
ETSformer
LightTS
Dlinear
FEDformer
Stationary
Autoformer
[12]
[20]
[5]
[16]
[15]
[19]
[18]
[21]
[11]
[17]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.343 0.367 0.329 0.367
0.404
0.426
0.364
0.387
0.338
0.375
0.375
0.398
0.374
0.400
0.345
0.372
0.379
0.419
0.386
0.398
0.505
0.475
192
0.381
0.383 0.367 0.385
0.450
0.451
0.398
0.404
0.374 0.387
0.408
0.410
0.400
0.407
0.380
0.389
0.426
0.441
0.459
0.444
0.553
0.496
336
0.414
0.404 0.399 0.410
0.532
0.515
0.428
0.425
0.410 0.411
0.435
0.428
0.438
0.438
0.413
0.413
0.445
0.459
0.495
0.464
0.621
0.537
720
0.475 0.438 0.454 0.439
0.666
0.589
0.487
0.461
0.478
0.450
0.499
0.462
0.527
0.502
0.474
0.453
0.543
0.490
0.585
0.516
0.671
0.561
Avg
0.403 0.398 0.387 0.400
0.513
0.496
0.419
0.419
0.400 0.406
0.429
0.425
0.435
0.437
0.403
0.407
0.448
0.452
0.481
0.456
0.588
0.517
ETTm2
96
0.181 0.264 0.175 0.259
0.287
0.366
0.207
0.305
0.187 0.267
0.189
0.280
0.209
0.308
0.193
0.292
0.203
0.287
0.192
0.274
0.255
0.339
192
0.246 0.304 0.241 0.302
0.414
0.492
0.290
0.364
0.249 0.309
0.253
0.319
0.311
0.382
0.284
0.362
0.269
0.328
0.280
0.339
0.281
0.340
336
0.307 0.343 0.305 0.343
0.597
0.542
0.377
0.422
0.321
0.351 0.314
0.357
0.442
0.466
0.369
0.427
0.325
0.366
0.334
0.361
0.339
0.372
720
0.407 0.398 0.402 0.400
1.730
1.042
0.558
0.524
0.408 0.403
0.414
0.413
0.675
0.587
0.554
0.522
0.421
0.415
0.417
0.413
0.433
0.432
Avg
0.286 0.327 0.281 0.326
0.757
0.610
0.358
0.404
0.291 0.333
0.293
0.342
0.409
0.436
0.350
0.401
0.305
0.349
0.306
0.347
0.327
0.371
ETTh1
96
0.383 0.391
0.414
0.419
0.423
0.448
0.479
0.464
0.384 0.402
0.494
0.479
0.424
0.432
0.386
0.400 0.376
0.419
0.513
0.491
0.449
0.459
192
0.433 0.419
0.460
0.445
0.471
0.474
0.525
0.492
0.436 0.429
0.538
0.504
0.475
0.462
0.437
0.432 0.420
0.448
0.534
0.504
0.500
0.482
336
0.475 0.441
0.501
0.466
0.570
0.546
0.565
0.515
0.491
0.469
0.574
0.521
0.518
0.488
0.481 0.459 0.459 0.465
0.588
0.535
0.521
0.496
720
0.475 0.464 0.500 0.488
0.653
0.621
0.594
0.558
0.521
0.500
0.562
0.535
0.547
0.533
0.519
0.516
0.506
0.507
0.643
0.616
0.514
0.512
Avg
0.441 0.429
0.469
0.454
0.529
0.522
0.541
0.507
0.458
0.450
0.542
0.510
0.491
0.479
0.456 0.452 0.440
0.460
0.570
0.537
0.496
0.487
ETTh2
96
0.288 0.337 0.302 0.348
0.745
0.584
0.400
0.440
0.340
0.374
0.340
0.391
0.397
0.437
0.333
0.387
0.358
0.397
0.476
0.458
0.346
0.388
192
0.375 0.392 0.388 0.400
0.877
0.656
0.528
0.509
0.402 0.414
0.430
0.439
0.520
0.504
0.477
0.476
0.429
0.439
0.512
0.493
0.456
0.452
336
0.423 0.434 0.426 0.433
1.043
0.731
0.643
0.571
0.452 0.452
0.485
0.479
0.626
0.559
0.594
0.541
0.496
0.487
0.552
0.551
0.482
0.486
720
0.434 0.446 0.431 0.446
1.104
0.763
0.874
0.679
0.462 0.468
0.500
0.497
0.863
0.672
0.831
0.657
0.463
0.474
0.562
0.560
0.515
0.511
Avg
0.380 0.402 0.387 0.407
0.942
0.684
0.611
0.550
0.414 0.427
0.439
0.452
0.602
0.543
0.559
0.515
0.437
0.449
0.526
0.516
0.450
0.459
Electricity
96
0.197
0.273 0.181 0.270
0.219
0.314
0.237
0.329
0.168 0.272
0.187
0.304
0.207
0.307
0.197
0.282
0.193
0.308
0.169
0.273
0.201
0.317
192
0.197
0.276 0.188 0.274
0.231
0.322
0.236
0.330
0.184
0.289
0.199
0.315
0.213
0.316
0.196
0.285
0.201
0.315
0.182
0.286
0.222
0.334
336
0.212
0.292 0.204 0.293
0.246
0.337
0.249
0.344
0.198 0.300
0.212
0.329
0.230
0.333
0.209
0.301
0.214
0.329
0.200
0.304
0.231
0.338
720
0.253
0.325
0.246
0.324
0.280
0.363
0.284
0.373
0.220 0.320 0.233
0.345
0.265
0.360
0.245
0.333
0.246
0.355
0.222 0.321
0.254
0.361
Avg
0.215
0.292 0.205 0.290
0.244
0.334
0.251
0.344
0.192 0.295
0.208
0.323
0.229
0.329
0.212
0.300
0.214
0.327
0.193
0.296
0.227
0.338
Exchange
96
0.082 0.198
0.177
0.218
0.158
0.230
0.202
0.261
0.107
0.234
0.085 0.204
0.116
0.262
0.088
0.218
0.148
0.278
0.111
0.237
0.197
0.323
192
0.176 0.297
0.225
0.259
0.206
0.277
0.242
0.298
0.226
0.344
0.182
0.303
0.215
0.359
0.176
0.315
0.271
0.380
0.219
0.335
0.300
0.369
336
0.328
0.412
0.278 0.297 0.272 0.335 0.287 0.335
0.367
0.448
0.348
0.428
0.377
0.466
0.313
0.427
0.460
0.500
0.421
0.476
0.509
0.524
720
0.861
0.700
0.354 0.348 0.398 0.418 0.351 0.386
0.964
0.746
1.025
0.774
0.831
0.699
0.839
0.695
1.195
0.841
1.092
0.769
1.447
0.941
Avg
0.362
0.402
0.259 0.281 0.259 0.315 0.271 0.320
0.416
0.443
0.410
0.427
0.385
0.447
0.354
0.414
0.519
0.500
0.461
0.454
0.613
0.539
Traffic
96
0.645
0.383
0.088 0.205 0.256
0.367
0.094 0.218
0.593
0.321
0.607
0.392
0.615
0.391
0.650
0.396
0.587
0.366
0.612
0.338
0.613
0.388
192
0.598
0.360
0.176 0.299 0.470
0.509
0.184 0.307
0.617
0.336
0.621
0.399
0.601
0.382
0.598
0.370
0.604
0.373
0.613
0.340
0.616
0.382
336
0.605
0.362
0.301
0.397
1.268
0.883
0.349
0.431
0.629
0.336
0.622
0.396
0.613
0.386
0.605
0.373
0.621
0.383
0.618
0.328
0.622
0.337
720
0.644
0.382
0.901
0.714
1.767
1.068
0.852
0.698
0.640 0.350 0.632
0.396
0.658
0.407
0.645
0.394
0.626
0.382
0.653
0.355
0.660
0.408
Avg
0.623
0.372 0.367
0.404
0.940
0.707
0.370
0.413
0.620
0.336
0.621
0.396
0.622
0.392
0.625
0.383
0.610
0.376
0.624
0.340
0.628
0.379
Weather
96
0.182 0.227
0.462
0.295
0.522
0.290
0.805
0.493
0.172 0.220
0.197
0.281
0.182
0.242
0.196
0.255
0.217
0.296
0.173 0.223
0.266
0.336
192
0.227 0.262
0.466
0.296
0.530
0.293
0.756
0.474
0.219 0.261
0.237
0.312
0.227
0.287
0.237
0.296
0.276
0.336
0.245
0.285
0.307
0.367
336
0.279 0.298
0.482
0.304
0.558
0.305
0.762
0.477
0.280
0.306
0.298
0.353
0.282
0.334
0.283
0.335
0.339
0.380
0.321
0.338
0.359
0.395
720
0.352
0.344
0.514
0.322
0.589
0.328
0.719
0.449
0.365
0.359
0.352 0.288
0.352
0.386
0.345
0.381
0.403
0.428
0.414
0.410
0.419
0.428
Avg
0.260 0.283
0.481
0.304
0.550
0.304
0.760
0.473
0.259 0.287
0.271
0.334
0.261
0.312
0.265
0.317
0.309
0.360
0.288
0.314
0.338
0.382
$1^{\text{st}}$ Count
8 20 16 13
1
0
1
0
7 5
0
1
0
0
1
0
5
0
1
1
0
0
$2^{\text{nd}}$ Count 13 7 7 15
1
3
4
3
6 5
2
1
0
0
2
2
0
0
5
4
0
0
$3^{\text{rd}}$ Count
5 5
4
3
3
3
2
2
11 20
4
0
3
0
6 4
2
1
0
1
0
1
B.4 Full Ablation Study
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 6. In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 11.
$l_{C+T}\leqslant min(l_{C},l_{T})$ (11)
B.5 Full Hyperparameter Sensitivity
The GPU memory in the experimental environment is limited, so we conducted a 2-dimensional traversal of the sample space size $|S|\in\{0,1,2,3,4,5,6\}$ and learning rate $lr\in\{1*10^{-2},5*10^{-2},1*10^{-1}\}$ in Table 8. In the case of sufficient GPU memory, more hyperparameters can be traversed. Based on the complete results of hyperparameter sensitivity, it can be seen that TEFN is an insensitive model to hyperparameters, which means that random hyperparameters can achieve stable and superior prediction performance. We use variance to describe the fluctuation of hyperparameters in Table 3.
Table 8: Full results of hyperparameter sensitivity
$|S|$ $lr$ Target Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
0 $1*10^{-2}$ 96 0.197 0.273 0.403 0.407 0.288 0.337 0.344 0.367 0.182 0.264 0.086 0.204 0.645 0.384 0.186 0.230
192 0.197 0.276 0.433 0.420 0.375 0.392 0.385 0.385 0.247 0.305 0.178 0.299 0.598 0.360 0.232 0.266
336 0.212 0.292 0.478 0.443 0.423 0.434 0.414 0.404 0.310 0.344 0.328 0.412 0.605 0.362 0.283 0.301
720 0.254 0.326 0.476 0.465 0.452 0.458 0.478 0.438 0.410 0.400 0.886 0.711 0.644 0.382 0.357 0.348
$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.289 0.338 0.344 0.368 0.182 0.265 0.082 0.198 0.645 0.384 0.182 0.227
192 0.197 0.277 0.433 0.420 0.376 0.393 0.383 0.385 0.247 0.305 0.179 0.300 0.598 0.360 0.232 0.266
336 0.213 0.294 0.476 0.441 0.425 0.437 0.415 0.406 0.308 0.344 0.340 0.421 0.605 0.363 0.284 0.301
720 0.254 0.327 0.478 0.465 0.453 0.459 0.478 0.439 0.408 0.399 0.911 0.722 0.644 0.382 0.352 0.344
$1*10^{-1}$ 96 0.197 0.274 0.383 0.391 0.291 0.340 0.347 0.369 0.182 0.264 0.082 0.198 0.645 0.383 0.183 0.228
192 0.197 0.277 0.433 0.419 0.376 0.394 0.383 0.385 0.257 0.314 0.180 0.300 0.598 0.360 0.244 0.274
336 0.214 0.295 0.476 0.441 0.424 0.435 0.416 0.406 0.309 0.344 0.340 0.421 0.605 0.363 0.292 0.308
720 0.255 0.327 0.482 0.469 0.454 0.460 0.477 0.439 0.408 0.400 0.937 0.732 0.644 0.382 0.352 0.345
1 $1*10^{-2}$ 96 0.197 0.274 0.384 0.392 0.289 0.337 0.349 0.370 0.182 0.265 0.083 0.199 0.645 0.385 0.186 0.229
192 0.197 0.277 0.438 0.421 0.377 0.392 0.382 0.386 0.246 0.304 0.178 0.299 0.598 0.361 0.232 0.265
336 0.212 0.292 0.475 0.441 0.425 0.435 0.418 0.407 0.307 0.343 0.335 0.418 0.606 0.363 0.280 0.299
720 0.253 0.325 0.476 0.464 0.452 0.458 0.482 0.442 0.409 0.400 0.883 0.709 0.644 0.382 0.355 0.347
$5*10^{-2}$ 96 0.198 0.275 0.386 0.393 0.294 0.341 0.346 0.369 0.182 0.265 0.088 0.205 0.645 0.385 0.194 0.239
192 0.197 0.277 0.438 0.422 0.375 0.392 0.383 0.385 0.247 0.305 0.184 0.304 0.598 0.360 0.241 0.274
336 0.212 0.292 0.475 0.441 0.453 0.448 0.414 0.404 0.308 0.343 0.361 0.434 0.605 0.363 0.282 0.300
720 0.255 0.327 0.475 0.465 0.458 0.461 0.478 0.439 0.407 0.398 0.886 0.708 0.644 0.383 0.352 0.345
$1*10^{-1}$ 96 0.198 0.275 0.388 0.395 0.295 0.342 0.345 0.369 0.182 0.265 0.089 0.206 0.645 0.383 0.202 0.249
192 0.197 0.277 0.434 0.420 0.376 0.394 0.381 0.384 0.247 0.305 0.176 0.297 0.598 0.360 0.239 0.276
336 0.212 0.292 0.476 0.441 0.423 0.435 0.415 0.404 0.309 0.343 0.343 0.423 0.605 0.363 0.289 0.306
720 0.256 0.328 0.476 0.466 0.463 0.465 0.478 0.439 0.408 0.398 0.892 0.710 0.644 0.383 0.360 0.353
2 $1*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.290 0.338 0.347 0.369 0.183 0.266 0.083 0.200 0.645 0.383 0.187 0.230
192 0.197 0.276 0.434 0.420 0.375 0.393 0.386 0.386 0.247 0.305 0.177 0.298 0.598 0.360 0.232 0.266
336 0.212 0.292 0.476 0.441 0.423 0.435 0.414 0.405 0.311 0.346 0.348 0.427 0.605 0.362 0.284 0.301
720 0.253 0.325 0.476 0.466 0.453 0.459 0.478 0.439 0.408 0.399 0.898 0.719 0.644 0.382 0.355 0.347
$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.289 0.338 0.345 0.368 0.182 0.265 0.084 0.202 0.646 0.384 0.183 0.228
192 0.197 0.277 0.434 0.420 0.377 0.394 0.384 0.387 0.247 0.305 0.177 0.298 0.598 0.360 0.228 0.263
336 0.212 0.292 0.477 0.443 0.424 0.436 0.415 0.405 0.308 0.343 0.353 0.430 0.605 0.363 0.293 0.309
720 0.254 0.325 0.478 0.466 0.458 0.461 0.477 0.439 0.408 0.400 0.930 0.732 0.644 0.382 0.355 0.347
$1*10^{-1}$ 96 0.197 0.274 0.384 0.391 0.289 0.339 0.347 0.368 0.182 0.265 0.088 0.205 0.646 0.385 0.186 0.231
192 0.197 0.277 0.438 0.423 0.378 0.395 0.387 0.387 0.247 0.305 0.191 0.310 0.598 0.361 0.241 0.278
336 0.212 0.293 0.480 0.446 0.425 0.437 0.416 0.406 0.308 0.343 0.403 0.463 0.605 0.363 0.299 0.316
720 0.254 0.326 0.475 0.466 0.463 0.465 0.478 0.439 0.410 0.402 0.880 0.707 0.644 0.382 0.357 0.350
3 $1*10^{-2}$ 96 0.197 0.274 0.385 0.393 0.290 0.339 0.343 0.368 0.182 0.265 0.083 0.199 0.645 0.383 0.183 0.227
192 0.197 0.277 0.434 0.420 0.376 0.393 0.383 0.385 0.248 0.306 0.183 0.304 0.598 0.360 0.232 0.266
336 0.212 0.293 0.477 0.443 0.426 0.435 0.414 0.405 0.308 0.344 0.338 0.420 0.606 0.363 0.282 0.300
720 0.253 0.325 0.477 0.465 0.475 0.473 0.476 0.438 0.407 0.399 0.943 0.739 0.644 0.382 0.353 0.346
$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.290 0.337 0.344 0.368 0.183 0.266 0.083 0.200 0.645 0.383 0.189 0.236
192 0.198 0.277 0.435 0.421 0.382 0.398 0.384 0.385 0.258 0.316 0.200 0.317 0.598 0.360 0.234 0.269
336 0.213 0.293 0.476 0.441 0.426 0.437 0.415 0.406 0.308 0.344 0.353 0.429 0.606 0.363 0.285 0.303
720 0.254 0.325 0.489 0.474 0.460 0.463 0.478 0.439 0.412 0.402 1.031 0.775 0.644 0.383 0.357 0.350
$1*10^{-1}$ 96 0.197 0.274 0.384 0.391 0.291 0.339 0.346 0.369 0.184 0.267 0.084 0.201 0.645 0.384 0.196 0.245
192 0.198 0.277 0.437 0.423 0.381 0.398 0.385 0.386 0.248 0.306 0.176 0.297 0.598 0.360 0.254 0.290
336 0.212 0.292 0.475 0.441 0.430 0.440 0.416 0.406 0.328 0.361 0.362 0.435 0.607 0.363 0.288 0.308
720 0.254 0.325 0.492 0.477 0.434 0.446 0.478 0.440 0.418 0.405 1.046 0.782 0.644 0.383 0.357 0.350
4 $1*10^{-2}$ 96 0.197 0.274 0.393 0.399 0.288 0.338 0.344 0.368 0.185 0.268 0.084 0.201 0.645 0.383 0.182 0.227
192 0.197 0.276 0.434 0.420 0.376 0.393 0.381 0.385 0.247 0.305 0.179 0.300 0.598 0.360 0.227 0.262
336 0.212 0.292 0.478 0.444 0.424 0.436 0.415 0.405 0.307 0.343 0.338 0.420 0.606 0.363 0.279 0.298
720 0.255 0.326 0.476 0.465 0.457 0.462 0.477 0.439 0.409 0.399 0.889 0.713 0.644 0.383 0.353 0.345
$5*10^{-2}$ 96 0.197 0.274 0.383 0.392 0.292 0.340 0.345 0.368 0.181 0.264 0.085 0.203 0.645 0.384 0.203 0.254
192 0.198 0.278 0.433 0.420 0.377 0.394 0.385 0.386 0.248 0.307 0.179 0.300 0.598 0.360 0.234 0.270
336 0.213 0.293 0.486 0.449 0.425 0.435 0.417 0.407 0.312 0.347 0.346 0.427 0.607 0.363 0.294 0.314
720 0.257 0.330 0.480 0.467 0.462 0.464 0.479 0.440 0.408 0.399 0.861 0.700 0.644 0.383 0.359 0.353
$1*10^{-1}$ 96 0.198 0.274 0.384 0.392 0.293 0.342 0.358 0.376 0.183 0.266 0.085 0.203 0.646 0.384 0.211 0.262
192 0.197 0.277 0.434 0.421 0.397 0.411 0.399 0.395 0.249 0.307 0.180 0.300 0.598 0.360 0.255 0.292
336 0.215 0.295 0.489 0.452 0.473 0.467 0.420 0.407 0.309 0.344 0.377 0.447 0.606 0.363 0.305 0.324
720 0.255 0.328 0.485 0.470 0.454 0.460 0.480 0.441 0.414 0.404 0.923 0.727 0.644 0.383 0.375 0.364
5 $1*10^{-2}$ 96 0.198 0.275 0.384 0.392 0.291 0.341 0.343 0.367 0.182 0.265 0.083 0.201 0.646 0.385 0.184 0.229
192 0.198 0.277 0.434 0.421 0.379 0.396 0.381 0.385 0.247 0.305 0.187 0.306 0.600 0.362 0.227 0.262
336 0.212 0.293 0.476 0.442 0.434 0.443 0.416 0.406 0.308 0.343 0.340 0.421 0.606 0.364 0.280 0.299
720 0.254 0.326 0.479 0.466 0.457 0.462 0.479 0.440 0.409 0.401 0.957 0.744 0.645 0.384 0.354 0.348
$5*10^{-2}$ 96 0.198 0.274 0.383 0.391 0.300 0.348 0.349 0.372 0.183 0.267 0.085 0.204 0.645 0.386 0.204 0.249
192 0.198 0.278 0.433 0.419 0.378 0.395 0.382 0.383 0.248 0.306 0.177 0.299 0.599 0.362 0.251 0.288
336 0.212 0.292 0.475 0.441 0.428 0.439 0.418 0.407 0.313 0.348 0.363 0.437 0.607 0.365 0.301 0.320
720 0.265 0.339 0.478 0.467 0.504 0.488 0.478 0.441 0.409 0.400 0.986 0.755 0.645 0.384 0.357 0.351
$1*10^{-1}$ 96 0.198 0.275 0.386 0.393 0.316 0.360 0.375 0.388 0.197 0.280 0.084 0.201 0.647 0.388 0.209 0.260
192 0.199 0.279 0.434 0.420 0.418 0.425 0.393 0.390 0.251 0.310 0.178 0.299 0.600 0.362 0.280 0.307
336 0.241 0.330 0.476 0.442 0.454 0.457 0.435 0.417 0.326 0.359 0.541 0.541 0.606 0.363 0.293 0.314
720 0.278 0.355 0.479 0.469 0.505 0.494 0.505 0.456 0.421 0.410 1.052 0.780 0.645 0.385 0.375 0.366
6 $1*10^{-2}$ 96 0.199 0.276 0.384 0.392 0.290 0.340 0.347 0.369 0.183 0.265 0.095 0.214 0.649 0.389 0.194 0.240
192 0.199 0.279 0.437 0.422 0.379 0.396 0.383 0.385 0.247 0.305 0.191 0.311 0.601 0.365 0.230 0.266
336 0.213 0.294 0.478 0.444 0.430 0.440 0.417 0.406 0.308 0.343 0.372 0.443 0.609 0.367 0.285 0.303
720 0.257 0.329 0.481 0.466 0.469 0.468 0.475 0.438 0.411 0.402 0.915 0.724 0.648 0.387 0.355 0.348
$5*10^{-2}$ 96 0.198 0.274 0.383 0.391 0.298 0.345 0.379 0.387 0.185 0.268 0.094 0.215 0.648 0.386 0.204 0.250
192 0.198 0.277 0.436 0.421 0.416 0.422 0.388 0.386 0.248 0.307 0.181 0.303 0.599 0.361 0.252 0.289
336 0.213 0.294 0.475 0.442 0.433 0.439 0.426 0.412 0.310 0.346 0.439 0.487 0.606 0.363 0.288 0.308
720 0.254 0.326 0.487 0.473 0.512 0.491 0.482 0.441 0.414 0.405 0.935 0.733 0.645 0.384 0.360 0.355
$1*10^{-1}$ 96 0.198 0.275 0.384 0.392 0.341 0.380 0.396 0.401 0.223 0.305 0.088 0.206 0.647 0.386 0.214 0.263
192 0.234 0.320 0.439 0.423 0.427 0.430 0.478 0.441 0.260 0.319 0.182 0.304 0.599 0.361 0.253 0.289
336 0.258 0.335 0.481 0.444 0.435 0.442 0.507 0.459 0.325 0.359 0.372 0.445 0.607 0.364 0.307 0.325
720 0.255 0.326 0.479 0.467 0.473 0.479 0.507 0.457 0.427 0.414 1.055 0.783 0.645 0.384 0.376 0.368

### v2

@@
 Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting
-Abstract
-In real-world scenarios, time series forecasting often demands timeliness, making research on model backbones a perennially hot topic.
+Abstract.
+In practical scenarios, time series forecasting necessitates timeliness, especially when dealing with large datasets.
+Consequently, the exploration of model architectures remains a perennially trending topic in research.
 To meet these performance demands, we propose a novel backbone from the perspective of information fusion.
@@
 TEFN is not a model that achieves the ultimate in single aspect, but a model that balances performance, accuracy, stability, and interpretability.
-Code is available at this repository: https://github.com/ztxtech/Time-Evidence-Fusion-Network.
-1 Introduction
-The changes of things are accompanied by the passage of time, and time series appear in more and more application scenarios, such as health diagnosis, financial analysis, weather forecasting, and so on.
-Optimizing structure of backbone is an effective way to improve the performance of time series forecasting, where linear-based models such as Dlinear often outperform transformer-based models such as Crossformer [18, 20].
-Therefore, we hope to propose a model that balances performance, prediction accuracy, stability, and interpretability to provide support for highly reliable scenarios with time requirements.
-Evidence theory is a model of uncertain reasoning commonly used for decision-making in the fusion of multiple information [14, 6].
-Meanwhile, evidence theory is often appeared in reasoning framework based on symbolic logic, with high interpretability.
-Denoeux has implemented evidence neural networks for classification tasks [8].
-Therefore, we propose the Time Evidence Fusion Network (TEFN) for regression tasks.
-The strategy adopted by TEFN is to divide the time series into two perspectives: time and channels, treating each channel/time as an independent information source, and using an information fusion framework for prediction in Figure 1.
-Therefore, TEFN is a simple and naive model, without any technical modifications, and is a new backbone.
-Figure 1: The channel dimension and time dimension of multivariate time series can be considered as different sources of information.
+PVLDB Reference Format:
+PVLDB, 14(1): XXX-XXX, 2020.
+doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
+doi:XX.XX/XXX.XX
+PVLDB Artifact Availability:
+The source code, data, and/or other artifacts have been made available at https://github.com/ztxtech/Time-Evidence-Fusion-Network.
+1. Introduction
+The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains.
+From health diagnostics, where the progression of diseases is tracked over time, to financial analysis, where market trends are monitored for investment strategies, and weather forecasting, where atmospheric changes are predicted to safeguard communities, the importance of time series forecasting cannot be overstated.
+However, as the volume of data grows exponentially, time series forecasting faces significant performance challenges.
+The sheer scale of big data necessitates more efficient and faster algorithms to process and predict within acceptable time frames, which has become a pressing issue in the field.
+Figure 1.
+The channel dimension and time dimension of multivariate time series can be considered as different sources of information.
 Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
-In TEFN, we propose a novel module Basic Probability Assignment (BPA) Module.
-Based on the evidence theory of decision-making process, BPA maps different sources of information to distributions related to the target outcome.
-The target distribution is achieved by the fuzzy membership of different fuzzy sets [7].
-In image processing, convolution condenses the required features of the target [4].
-Unlike convolution, BPA is an expansion process in Figure 2.
-On the contrary, BPA considers all possibilities in the event space composed of subsets of the sample space, expanding on different events, which is more effective for simple structures like time series and can fully explore the different hidden information inside.
-For different distributions generated by multiple information sources, we consider sources as samples and use the expected forecasting values as the final result.
-Figure 2: Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
-In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability.
-At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects.
-TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20] or Patch [12].
-For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model.
-The contributions of this article are as follows:
-  • This article proposes a new evidential backbone for extracting potential distributions of simple data structures.
-  • A neural network TEFN based on BPA for time series forecasting has been proposed, which leads previous research in accuracy, efficiency, stability, and interpretability.
-  • This article explicitly introduces relevant concepts from evidence theory, such as mass function, into the structure of neural networks, providing a new implementation approach for uncertainty reasoning.
-2 Related Work
-2.1 Dempster-Shafer Theory
-Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [14, 6].
+In response to these challenges, researchers have been focusing on optimizing the structure of the backbone models used in time series forecasting.
+It has been observed that linear-based models, such as Dlinear (Zeng et al., 2023), often exhibit superior performance compared to transformer-based models like Crossformer (Zhang and Yan, 2023), particularly in terms of computational efficiency and speed.
+These findings highlight the potential of structural optimizations in enhancing the capabilities of time series forecasting models.
+Nevertheless, the quest for an ideal model that balances performance, prediction accuracy, stability, and interpretability continues.
+The aim is to develop a model that can provide robust support for high-stakes scenarios where timely and reliable predictions are critical, thereby bridging the gap between theoretical advancements and practical application needs.
+Evidence theory, also known as Dempster-Shafer theory, provides a mathematical framework for representing and combining degrees of belief in the presence of uncertainty (Shafer, 1976; DEMPSTER, 1967).
+It is particularly useful in situations where information is incomplete, ambiguous, or comes from diverse sources.
+The theory is grounded in the concept of a belief function, which quantifies the degree to which a hypothesis is supported by the available evidence, without necessarily committing to a probability distribution over the possible outcomes.
+One of the key strengths of evidence theory is its ability to handle the fusion of multiple information sources effectively.
+This makes it a powerful tool in fields such as data fusion, sensor networks, and multi-criteria decision-making, where diverse pieces of information need to be integrated to form a coherent union.
+In the realm of symbolic logic, evidence theory often appears as a reasoning framework that offers high interpretability.
+Unlike traditional probability theory, which requires precise probability assignments, evidence theory allows for a more nuanced representation of uncertainty.
+It does this through the use of belief and plausibility measures, which provide a range of values for the likelihood of an event occurring, rather than a single probability.
+An innovative application of evidence theory in machine learning is the development of evidence neural networks by Denoeux (Denoeux, 2000).
+These networks are designed for classification tasks and integrate the principles of evidence theory into the architecture of neural networks.
+By doing so, they enhance the network’s ability to handle uncertainty and provide a more robust and interpretable decision-making process.
+The evidence neural networks output belief masses that can be directly interpreted as measures of evidence supporting each class, thus offering a clear understanding of the network’s reasoning process.
+Therefore, we introduce the Time Evidence Fusion Network (TEFN), a novel architecture designed specifically for regression tasks within the realm of time series forecasting.
+The conceptual underpinning of TEFN is the innovative strategy of dissecting the time series data into dual perspectives—time and channels—thereby treating each channel and temporal instance as a distinct and autonomous source of information.
+This approach allows us to leverage the principles of evidence theory to fuse these diverse information streams within an integrated framework, aiming to enhance the accuracy and reliability of predictions as depicted in Figure 1.
+In essence, TEFN is built upon the premise of simplicity and naivety in its design philosophy.
+It refrains from incorporating complex technical modifications that often render models opaque and difficult to interpret.
+Instead, TEFN maintains a straightforward structure that prioritizes clarity and ease of understanding, making it a transparent and accessible backbone for time series analysis.
+By adopting this minimalist approach, TEFN emerges as a refreshing alternative in the field, offering a new foundation upon which more sophisticated forecasting models can potentially be built, while retaining the interpretability and trustworthiness that are hallmarks of evidence theory.
+In the architecture of TEFN, we introduce a pivotal component termed the Basic Probability Assignment (BPA) Module, which stands as the core of our evidence-theoretic approach to time series forecasting.
+The BPA Module operates on the fundamental principles of evidence theory, meticulously mapping various sources of information into mass distributions that are intrinsically linked to the desired target outcome.
+This mapping is achieved through the application of fuzzy membership functions, which assign membership grades to different fuzzy sets, thereby constructing a comprehensive target distribution as proposed by Deng (Deng, 2015).
+In contrast to the convolutional operations prevalent in image processing, which selectively condense features pertinent to the target through filters (Bengio et al., 1993), the BPA Module embodies an expansionary process, as illustrated in Figure 2.
+Rather than narrowing down to specific features, BPA embraces an expansive perspective, contemplating all potentialities within the event space, which is a collection of subsets derived from the sample space.
+This expansion across various events is particularly advantageous for the simplistic structures inherent in time series data, enabling a thorough exploration of the myriad hidden informational layers within.
+For the diverse distributions yielded by the multitude of information sources, we treat each source as an individual sample, and by amalgamating these samples, we utilize the expected forecasting values as the ultimate prediction, thereby delivering a robust and informed forecast that encapsulates the collective wisdom of all contributing sources.
+Figure 2. Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
+In this paper, we delve into the practical nuances of the Time Evidence Fusion Network (TEFN), examining its precision, memory consumption, training velocity, sensitivity to parameters, and interpretability.
+To maintain equitable comparisons, all experimental data and configuration files are sourced from open-source initiatives.
+TEFN has partially achieved state-of-the-art (SOTA) performance on various real-world datasets, with a notable reduction in training duration and model complexity when contrasted with Transformer-based models such as Crossformer (Zhang and Yan, 2023) and Patch-based models like PatchTST (Nie et al., 2023).
+In terms of random hyperparameter selection, TEFN demonstrates minimal variability, affirming its stability.
+The contributions of this manuscript are summarized as follows:
+  • The proposed method is well-suited for large datasets, showcasing exceptional performance.
+  • A novel evidential backbone BPA designed for extracting potential distributions from simple data structures is introduced.
+  • A BPA-based neural network, TEFN, is introduced for time series forecasting, outpacing prior research in terms of accuracy, efficiency, stability, and interpretability.
+  • The manuscript elucidates the incorporation of evidence theory concepts, such as the mass function, into the architecture of neural networks, offering a novel approach to uncertainty reasoning in machine learning.
+This paper is organized as follows.
+We begin by reviewing relevant literature in the Section 2.
+Next, in the Section 3, we delve into the architecture and underlying mathematical principles of TEFN.
+In the Section 4, we present a series of experiments that demonstrate TEFN’s superior performance.
+Finally, we conclude the paper by summarizing our key findings in the Section 5.
+2. Related Work
+2.1. Dempster-Shafer Theory
+Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory (Shafer, 1976; DEMPSTER, 1967).
 It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities.
 When combining two mass functions, denoted as $m_{1}$ and $m_{2}$, the resulting fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation 1.
-$m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
-2.2 Evidence Decision Making
-The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [7].
+The symbols $A$, $B$, and $C$ represent focal elements, which are typically combinations of all independent samples (or target labels, such as the classification targets).
+(1) $m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$
+2.2. Basic Probability Assignment
+BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions.
+Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making.
+The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation 2 where $w$ and $b$ is weight and bias of the membership function.
+BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation 3.
+(2) $\begin{split}m=\mu(x)=wx+b\end{split}$
+(3) $\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$
+The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information.
+2.3. Evidence Decision Making
+The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) (Deng, 2015; Zhan et al., 2024).
 BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution.
 Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution.
+The most classic fusion method is the DSR (Shafer, 1976; DEMPSTER, 1967) mentioned in Section 2.1.
 Finally, decision making transpires as the progression from the attained distribution towards the designated target.
-3 Time Evidence Fusion Network
-The structure of TEFN is naive, and well performance can be achieved through a structure with only $2$ BPA module similar to only $2$ convolution kernels.
-We believe that the new architecture of TEFN is more like a pure backbone model.
-The overall structure of TEFN is in Figure 3.
-Figure 3: The overall structure of TEFN
-3.1 Time Normalization and De-normalization
-In order to accelerate converging to local optima more quickly, we referred to the normalization methods of Stationary in TEFN [11].
-The normalization process calculates the mean $\mu$ in Equation 2 and variance difference $\sigma^{2}$ in Equation 3 for each piece of the time series $x$.
-The normalization and de-normalization processes are a pair of symmetric operations in Equation 4 and Equation 5, where $Net(\cdot)$ is the neural network operation performed on the normalized time series $x_{Norm}$.
-$\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$ (2)
-$\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$ (3)
-$x_{Norm}=\frac{x-\mu}{\sigma}$ (4)
-$x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (5)
-3.2 Time Dimension Projection
-In TEFN, a linear layer is used to project the length $L_{in}$ of the time dimension to length $L_{in}+L_{pred}$ where $L_{pred}$ is prediction length of current target.
-3.3 Basic Probability Assignment Module
-Figure 4: The operation process of a BPA module for a time series $T$ of length $L=5$.
+The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignostic probability transformation (PPT) (Dubois et al., 1993) in Equation 4, where $x$ is a target label and $e$ are combinations of all possible target labels.
+(4) $p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$
+Figure 3. The overall structure of TEFN
+2.4. Linear Time Series Model
+There are many linear time series forecasting models (LTSFM).
+LTSFM offer several advantages for modeling and forecasting time series data.
+Firstly, their simplicity and transparency make them easy to implement, interpret, and analyze.
+The parameters of linear models directly represent the relationships between input features and the output, providing insights into the underlying dynamics of the time series.
+Secondly, LTSFM exhibit high efficiency, requiring fewer computational resources and hyperparameters compared to complex models like transformer-based models (Vaswani, 2017; Zhang and Yan, 2023).
+This leads to faster training and inference times, making them suitable for real-time applications and large datasets.
+A classic LTSFM model is Dlinear (Zeng et al., 2023).
+DLinear serves as a simple yet effective baseline for long-term time series forecasting.
+It is a model that employs a seasonal-trend decomposition technique to separate the trend and seasonal components of the time series data.
+Each component is then independently modeled using a single-layer linear regression.
+The final prediction is obtained by summing the outputs of the two linear regressions.
+TEFN shares similarities with Dlinear in that both models utilize linear regression for prediction. However, there are key differences:
+  • DLinear employs a simpler decomposition method compared to TEFN, which may limit its ability to capture intricate temporal dynamics.
+  • DLinear focuses on capturing linear relationships and trends in the data, while TEFN focuses on the mining of internal hidden relationships, so it strengthens predictions through dimensionality increase and uncertain reasoning.
+  • TEFN is a general framework that does not require identification and prediction of trends and periodicity, while Dlinear is more like a feature modeling prediction of time series.
+3. Time Evidence Fusion Network
+The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels.
+We posit that the novel architecture of TEFN embodies the essence of a pure backbone model, stripped of domain-specific intricacies.
+This minimalist approach allows TEFN to focus solely on capturing the fundamental temporal patterns within the data, making it a versatile and adaptable foundation for various time series analysis tasks.
+The overall structure of TEFN is depicted in Figure 3.
+Figure 4.
+The operation process of a BPA module for a time series $T$ of length $L=5$.
 The sample space size in the BPA module is set as $|S|=2$, and the event space size is set as $|2^{S}|=2^{|S|}=4$.
 The operation of the BPA module is similar to that of the convolution module, but the BPA module diffuses different masses instead of summing them in the convolution.
-According to Evidence Theory, each dimension of a time series can be reduced to a mass function in the same set of event spaces.
-The event space of evidence theory usually chooses power sets $2^{S}$ of sample space $S$.
-If it is a classification problem, the sample space is all the labels.
-For regression problems, the sample space needs to be treated as anonymous samples, where the samples are mutually exclusive.
-Similarly, whether it is a classification or regression problem, each dimension of the data will generate a mass distribution in the event space.
-Fuzzy numbers commonly used in the generation of mass distribution of evidence.
-Here we use a parameterized fuzzy membership $\mu$ in Equation 6, where $D$ (time dimension $T$ and channel dimension $C$ are available) is the dimension that generates the mass function, $j$ is corresponding to dimension $D$ and $i$ is corresponding to the other one.
-Parameter $w$ and $b$ represent the slope and intercept of the membership function, and $k$ is corresponding to event dimension $F$.
-$m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (6)
-3.4 Expectation Fusion
-For multivariate time series in shape $|T|*|C|$, there are two dimensions: time dimension $T$ and channel dimension $C$.
-So, the different mass distributions $m_{C},m_{T}$ can be generated simultaneously.
-And the fusion method here uses expectations in Equation 7 where $y_{j,k}$ is parameter and $m^{\prime}_{i,j,k}$ is corresponding mass function.
-Summing part $y_{j,k}*m^{\prime}_{i,j,k}$ can be regarded as a linear transform of $m_{i,j,k}$, and expectation is calculated by summing $m_{i,j,k}$ in Equation 7.
-There are two reasons why DSR are not used:
-  1. 1.
-    The mathematical operation corresponding to DSR is the point-to-point multiplication of mass distributions within the same dimension, which has a high computational complexity.
-At the same time, if the dimension is too high, it will lead to the disappearance of accuracy within a finite precision.
-  2. 2.
-    DSR is sensitive to one extreme distribution, such as a distribution with only one element being $1$, which can result in the final fusion result of the entire mass being directly equal to that distribution.
-$\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$ (7)
-At the end, results of time dimension $T$ and channel dimension $C$ will aggregate in Equation 8 where $\hat{y}_{norm}$ is normalized predicted results.
-The final prediction results are restored through the de-normalization layer.
-$\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (8)
-4 Experiments
-Dataset
-We will use TEFN for long-term prediction tasks.
-Table 1 shows the basic settings used in the experiment, including the datasets.
-In terms of datasets, we have selected $5$ common time series datasets, including Electricity [1], ETT (4 subsets) [10], Exchange [17], Traffic [2], Weather [3].
-More dataset features are in the Appendix Table 5.
-Table 1: Datasets and Basic Settings
-Tasks
-Datasets
-Metrics
-Series Length
-Long-term
-ETT (4 subsets), Electricity,
-MSE, MAE 96$\sim$720
-Forecasting
-Traffic, Weather, Exchange
-Baseline
-The comparative experimental models include PatchTST (ICLR 2023) [12], Crossformer (ICLR 2023) [20], TiDE [5], TimeNet (ICLR 2023) [16], ETSformer [15], LightTS [19], Dlinear (AAAI 2023) [18], FEDformer (ICML 2022) [21], Stationary (NeurIPS 2022) [11] and Autoformer (NeurIPS 2021) [17] .
-The configurations of the comparative experiments are from the Time Series Library (https://github.com/thuml/Time-Series-Library).
-More implementation details are in the Appendix Section A.
-4.1 Forecasting Result
-TEFN partially achieved SOTA performance in long-term forecasting target in Table 2.
-It is worth noting that in the case of a simple structure with tiny parameters in Section 4.4, achieved approximate or even better performance than previous SOTA models with larger parameters, such as PatchTST [12].
-Full forecasting results are in the Appendix Table 7.
-Table 2: Prediction error. All results are the average of four predicted lengths $\{96,192,336,720\}$.
+3.1. Time Normalization and De-normalization
+To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model (Liu et al., 2022).
+This process computes the mean $\mu$ as shown in Equation 5 and variance $\sigma^{2}$ as shown in Equation 6 for each segment of the time series $x$.
+The subsequent normalization and de-normalization operations, presented in Equation 7 and Equation 8, respectively, act as symmetrical transformations that standardize the time series data.
+The neural network operation $Net(\cdot)$ is then applied to the normalized time series $x_{Norm}$, ensuring that the model learns from data with consistent scale and distribution.
+This normalization strategy effectively mitigates the influence of outliers and initial data scale, fostering faster convergence and improved stability during training.
+(5) $\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$
+(6) $\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$
+(7) $x_{Norm}=\frac{x-\mu}{\sigma}$
+(8) $x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$
+3.2. Time Dimension Projection
+In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation 9.
+Here, $L_{pred}$ represents the desired length of the prediction horizon.
+This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet (Wu et al., 2023).
+(9) $x^{\prime}=project(x)=x\times W_{p}+b$
+By extending the input vector to match the prediction length, the time projection layer facilitates the direct generation of future values based on the learned temporal patterns.
+This approach simplifies the model design and enables efficient prediction without the need for iterative or recursive computations, as commonly observed in autoregressive models.
+3.3. Basic Probability Assignment Module
+In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data.
+According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space.
+This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data.
+For classification problems, $S$ corresponds to all the possible labels, while for regression problems, $S$ is treated as a set of anonymous samples, ensuring mutual exclusivity among them.
+This representation allows us to capture the fuzzy nature of the data, where each dimension generates a mass distribution within the event space.
+To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation 10.
+Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$.
+The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$.
+The fuzzy membership function $\mu_{k}$ is defined in Equation 10 where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function, respectively.
+These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space.
+(10) $m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$
+By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions.
+This approach provides a principled way to handle the inherent fuzziness and ambiguity in real-world time series, leading to improved performance and reliability in various applications.
+To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module.
+Instead of the linear form presented in Equation 10, we can adopt non-linear functions that capture more complex relationships within the data.
+Several options exist for non-linear membership functions, including:
+  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation 3, capturing the likelihood of each data point belonging to different membership levels.
+The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
+  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation 11.
+The sigmoid function can capture both gradual and abrupt changes in membership values.
+    (11) $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$
+  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation 12 where $g(x)$ is a non-linear function.
+This approach is useful for capturing complex patterns and relationships within the data.
+    (12) $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$
+By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data.
+This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
+3.4. Expectation Fusion
+TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$.
+This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$.
+To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation 13.
+Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
+(13) $\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$
+The fusion operation in Equation 13 can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions.
+The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
+TEFN deliberately avoids using the Dempster-Shafer Rule of Combination (DSR) for several reasons.
+Firstly, the computational complexity of DSR is significantly higher compared to the expectation-based fusion method.
+Additionally, DSR involves point-to-point multiplication of mass distributions within the same dimension, which can lead to accuracy loss when dealing with high-dimensional data due to finite precision limitations.
+Secondly, DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1.
+This sensitivity can result in the final fused result being dominated by that single distribution, potentially leading to conflicts and loss of information from other distributions.
+By employing the expectation-based fusion method, TEFN achieves a balance between computational efficiency and accuracy.
+This approach effectively integrates information from different dimensions while mitigating the risks of conflicts and accuracy loss associated with DSR.
+The final normalized prediction results are then aggregated across dimensions using the equation in Equation 14 and restored to their original scale through the de-normalization layer.
+(14) $\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$
+4. Experiments
+The experimental section aims to comprehensively evaluate TEFN’s performance in long-term time series forecasting.
+Utilizing diverse datasets, TEFN will be benchmarked against established models, measuring accuracy, efficiency, and robustness.
+We will analyze its performance across different prediction horizons, assess its sensitivity to hyperparameters, and conduct ablation studies to understand the impact of individual components.
+Lastly, we will explore TEFN’s interpretability by analyzing the basic probability assignments generated by its BPA module.
+4.1. Dataset
+Our research employs the TEFN for the challenging task of long-term time series prediction.
+To ensure the model’s effectiveness and robustness across diverse scenarios, we meticulously selected five representative and extensively used datasets.
+These datasets encompass a wide range of real-world applications, providing a comprehensive evaluation of TEFN’s performance.
+The selected datasets are:
+  • Electricity (ecl, [n.d.]): This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time.
+It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
+  • ETT (4 subsets) (Li et al., 2021): This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges.
+It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
+  • Exchange (Wu et al., 2021): This dataset captures exchange rate fluctuations between major currencies.
+It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
+  • Traffic (tra, [n.d.]): This dataset contains traffic flow data collected from sensors in a city.
+It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
+  • Weather (wea, [n.d.]): This dataset includes historical weather data, encompassing various meteorological parameters.
+It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
+Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities.
+The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability.
+More dataset features are in the Table 1.
+Table 1. The quantitative features of the dataset
+Dataset Dimension Dataset Size
+Electricity 321 (18317, 2633, 5261)
+ETTm1, ETTm2 7 (34465, 11521, 11521)
+ETTh1, ETTh2 7 (8545, 2881, 2881)
+Traffic 862 (12185, 1757, 3509)
+Weather 21 (36792, 5271, 10540)
+Exchange 8 (5120, 665, 1422)
+Table 2.
+This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks.
+The models are evaluated on four different prediction lengths: $L_{pred}\in\{96,192,336,720\}$, providing a thorough assessment of their performance under varying forecasting horizons.
+The color coding in the table highlights the ranking of each model’s performance across all datasets and prediction lengths, allowing for easy identification of the top-performing models.
 Models
 TEFN
+iTransformer
+Rlinear
 PatchTST
@@
 Autoformer
-[12]
-[20]
-[5]
-[16]
-[15]
-[19]
-[18]
-[21]
-[11]
-[17]
 Metric
@@
 MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-MSE
-MAE
-ETTm1
-0.403 0.398 0.387 0.400
-0.513
-0.496
-0.419
-0.419
-0.400 0.406
-0.429
-0.425
-0.435
-0.437
-0.403
-0.407
-0.448
-0.452
-0.481
-0.456
-0.588
-0.517
-ETTm2
-0.286 0.327 0.281 0.326
-0.757
-0.610
-0.358
-0.404
-0.291 0.333
-0.293
-0.342
-0.409
-0.436
-0.350
-0.401
-0.305
-0.349
-0.306
-0.347
-0.327
-0.371
-ETTh1
-0.441 0.429
-0.469
-0.454
-0.529
-0.522
-0.541
-0.507
-0.458
-0.450
-0.542
-0.510
-0.491
-0.479
-0.456 0.452 0.440
-0.460
-0.570
-0.537
-0.496
-0.487
-ETTh2
-0.380 0.402 0.387 0.407
-0.942
-0.684
-0.611
-0.550
-0.414 0.427
-0.439
-0.452
-0.602
-0.543
-0.559
-0.515
-0.437
-0.449
-0.526
-0.516
-0.450
-0.459
-Electricity
-0.215
-0.292 0.205 0.290
-0.244
-0.334
-0.251
-0.344
-0.192 0.295
-0.208
-0.323
-0.229
-0.329
-0.212
-0.300
-0.214
-0.327
-0.193
-0.296
-0.227
-0.338
-Exchange
-0.362
-0.402
-0.259 0.281 0.259 0.315 0.271 0.320
-0.416
-0.443
-0.410
-0.427
-0.385
-0.447
-0.354
-0.414
-0.519
-0.500
-0.461
-0.454
-0.613
-0.539
-Traffic
-0.623
-0.372 0.367
-0.404
-0.940
-0.707
-0.370
-0.413
-0.620
-0.336
-0.621
-0.396
-0.622
-0.392
-0.625
-0.383
-0.610
-0.376
-0.624
-0.340
-0.628
-0.379
-Weather
-0.260 0.283
-0.481
-0.304
-0.550
-0.304
-0.760
-0.473
-0.259 0.287
-0.271
-0.334
-0.261
-0.312
-0.265
-0.317
-0.309
-0.360
-0.288
-0.314
-0.338
-0.382
-$1^{\text{st}}$ Count
-8 20 16 13
-1
-0
-1
-0
-7 5
-0
-1
-0
-0
-1
-0
-5
-0
-1
-1
-0
-0
-$2^{\text{nd}}$ Count 13 7 7 15
-1
-3
-4
-3
-6 5
-2
-1
-0
-0
-2
-2
-0
-0
-5
-4
-0
-0
-$3^{\text{rd}}$ Count
-5 5
-4
-3
-3
-3
-2
-2
-11 20
-4
-0
-3
-0
-6 4
-2
-1
-0
-1
-0
-1
-4.2 Hyperparameter Sensitivity
-To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$.
-We take the Electricity-96 task as an example.
-TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 5.
-It is worth noting that the order of magnitude of the MSE in Figure 5 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small.
-Full hyperparameter sensitivity experiment results are in the Appendix Table 8.
-(a) $lr$ $|S|$MSE
-(b) Violin diagram of $|S|$ and MSE
-Figure 5: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
-Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table 3.
-The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation.
-Random parameter selection within the traversal range of TEFN yields stable prediction results.
-Table 3: The variance of model error for traversal learning rate and sample space overhaul
-Metric Dataset
-($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
-MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
-192 0.0001 0.0000 0.0003 0.0004 0.0000 0.0000 0.0000 0.0002
-336 0.0001 0.0000 0.0002 0.0004 0.0000 0.0023 0.0000 0.0001
-720 0.0000 0.0000 0.0004 0.0001 0.0000 0.0038 0.0000 0.0001
-MAE 96 0.0000 0.0000 0.0001 0.0001 0.0001 0.0000 0.0000 0.0002
-192 0.0001 0.0000 0.0001 0.0001 0.0000 0.0000 0.0000 0.0002
-336 0.0001 0.0000 0.0001 0.0001 0.0000 0.0009 0.0000 0.0001
-720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
-4.3 Ablation Study
-TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$.
-In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table 4.
-Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error.
-Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (error of Electricty only varies $0.001$).
-It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
-Table 4: The results of the ablation experiment
-Metric Models
-Dataset
-Electricity
-ETTh1
-ETTh2
-ETTm1
-ETTm2
-Exchange
-Traffic
-Weather
-MSE C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
-C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
-T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
-MAE C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
-C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
-T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
-$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
-4.4 Efficiency Comparison
-Figure 6: Model efficiency comparison.
-Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
-TEFN is a compact model compared to the parameters of the Transformer.
-In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison.
-In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample.
-Different models occupy varying amounts of GPU and memory.
-Therefore, we used the size of the saved binary model files as the model size.
-We visualized this with a bubble chart in Figure 6 .
-It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters.
-Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency.
-TEFN maintains almost Dlinear [18] performance and almost PatchTST [12] accuracy, making it a performance error balanced model.
-4.5 Interpretability Analysis
-TEFN will allocate mass based on the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, and mass indicates the confidence reflecting support of different possible events.
-This support allocation is interpretable because BPA allocation utilizes fuzzy membership $\mu(x)=wx+b$ in Figure 7.
-Essentially, different fuzzy sets are constructed through TEFN for inference.
-(a) Channel BPA
-(b) Time BPA
-Figure 7: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time.
-The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
-4.6 Limitation and Future Prospects
-TEFN is a naive and unmodified model.
-We lack skills in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method.
-Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well.
-In addition, for the fusion of mass distribution in the face of the failure of the classic Dempster Shafer rule for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern.
-In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels.
-In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution.
-Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic.
-Effective parameter initialization may exist, for example, initializing corresponding parameters in a non neural network way may lead to faster convergence.
-5 Conclusion
-This paper treats different channels and different time points of time series as multiple sources of information.
-By utilizing evidence theory to construct different basic probability assignments, TEFN discards irrelevant time points and channels.
-Experimentally, TEFN is a parameter-insensitive and stable prediction model with a small parameter count and fast training speed.
-In future work, we will explore how to optimize basic probability assignments and information fusion in neural networks.
-Acknowledgments
-The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078).
-Appendix A Implementation Details
-Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU.
-Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores , 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB.
-The proposed model TEFN is implemented in the PyTorch [13].
-For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations.
-The loss function selected during the experiment was MSE, and the optimizer selected Adam [9].
-We used Mean Absolute Error (MAE) and Mean Square Error (MSE) in Equation 9 and Equation 10 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
-$MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (9)
-$MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (10)
-Appendix B Supplementary Experimental Results
-B.1 Dataset Feature
-In the experimental section, we mainly referred to the experimental dataset and experimental settings of TimesNet.
-The experimental settings of TimesNet are comprehensive and reasonable [16].
-Table 5 shows the quantitative features of the dataset used in the experiment.
-Dataset size corresponds to the training set, validation set, and test set.
-Each dataset has four forecasting tasks of different lengths $L\in\{96,192,336,720\}$.
-Table 5: The quantitative features of the dataset
-Dataset Dimension Dataset Size
-Electricity 321 (18317, 2633, 5261)
-ETTm1, ETTm2 7 (34465, 11521, 11521)
-ETTh1, ETTh2 7 (8545, 2881, 2881)
-Traffic 862 (12185, 1757, 3509)
-Weather 21 (36792, 5271, 10540)
-Exchange 8 (5120, 665, 1422)
-B.2 Forecasting Visualization
-We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models.
-We take the task of Electricity-96 [1] as an example and selected typical models PatchTST [12], Crossformer [20], TimesNet [16], Dlinear [18] and Stationary [11] for visualization in Figure 8.
-From the visualization results, TEFN fits the details and trends of time series better.
-(a) TEFN (Ours)
-(b) PatchTST
-(c) Crossformer
-(d) TimesNet
-(e) Dlinear
-(f) Stationary
-Figure 8: Visualization of Electricity-96 tasks
-B.3 Full Comparative Results
-In the comparative experiment, we conducted a thorough comparison of 5 datasets and 10 comparative models in terms of forecasting length $l\in\{96,192,336,720\}$ in Table 7.
-It can be seen that TEFN has positive predictive performance on the experiment dataset, achieving SOTA in multiple forecasting tasks across multiple datasets.
-Table 6: Full results of the ablation experiment
-Metric Target Model Dataset
-Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
-MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
-C 0.196$\downarrow$ 0.386$\uparrow$ 0.288 0.339$\downarrow$ 0.185$\uparrow$ 0.083$\uparrow$ 0.644$\downarrow$ 0.185$\uparrow$
-T 0.197 0.386$\uparrow$ 0.291$\uparrow$ 0.356$\uparrow$ 0.182$\uparrow$ 0.083$\uparrow$ 0.645 0.195$\uparrow$
-192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
-C 0.196$\downarrow$ 0.435$\uparrow$ 0.375 0.382$\uparrow$ 0.246 0.179$\uparrow$ 0.597$\downarrow$ 0.230$\uparrow$
-T 0.197 0.435$\uparrow$ 0.375 0.390$\uparrow$ 0.247$\uparrow$ 0.175$\downarrow$ 0.598 0.240$\uparrow$
-336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
-C 0.210$\downarrow$ 0.479$\uparrow$ 0.423 0.411$\downarrow$ 0.307 0.336$\uparrow$ 0.604$\downarrow$ 0.280$\uparrow$
-T 0.212 0.479$\uparrow$ 0.414$\downarrow$ 0.423$\uparrow$ 0.307 0.324$\downarrow$ 0.606$\uparrow$ 0.291$\uparrow$
-720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
-C 0.253 0.478$\uparrow$ 0.457$\uparrow$ 0.472$\downarrow$ 0.407 0.947$\uparrow$ 0.642$\downarrow$ 0.354$\uparrow$
-T 0.254$\uparrow$ 0.479$\uparrow$ 0.442$\uparrow$ 0.487$\uparrow$ 0.408$\uparrow$ 0.871$\uparrow$ 0.643$\downarrow$ 0.364$\uparrow$
-Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
-C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
-T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
-MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
-C 0.273$\downarrow$ 0.394$\uparrow$ 0.339$\uparrow$ 0.366$\downarrow$ 0.269$\uparrow$ 0.199$\uparrow$ 0.384 0.228$\uparrow$
-T 0.274 0.393$\uparrow$ 0.339$\uparrow$ 0.377$\uparrow$ 0.265$\uparrow$ 0.199$\uparrow$ 0.383$\downarrow$ 0.234$\uparrow$
-192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
-C 0.275$\downarrow$ 0.421$\uparrow$ 0.393$\uparrow$ 0.385$\uparrow$ 0.305$\uparrow$ 0.300$\uparrow$ 0.361$\uparrow$ 0.264$\uparrow$
-T 0.276 0.421$\uparrow$ 0.390$\downarrow$ 0.392$\uparrow$ 0.305$\uparrow$ 0.296$\downarrow$ 0.361$\uparrow$ 0.270$\uparrow$
-336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
-C 0.291$\downarrow$ 0.443$\uparrow$ 0.434$\downarrow$ 0.403$\downarrow$ 0.343 0.419$\uparrow$ 0.363 0.298
-T 0.292 0.445$\uparrow$ 0.425$\downarrow$ 0.413$\uparrow$ 0.342$\downarrow$ 0.410$\downarrow$ 0.362$\downarrow$ 0.306$\uparrow$
-720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
-C 0.325 0.463$\downarrow$ 0.462$\uparrow$ 0.436$\downarrow$ 0.399 0.739$\uparrow$ 0.382 0.347$\uparrow$
-T 0.324$\downarrow$ 0.469$\uparrow$ 0.454$\uparrow$ 0.449$\uparrow$ 0.398$\downarrow$ 0.702$\uparrow$ 0.382 0.354$\uparrow$
-Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
-C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
-T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
-$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
-Table 7: Full results of comparative results
-Models
-TEFN
-PatchTST
-Crossformer
-TiDE
-TimesNet
-ETSformer
-LightTS
-Dlinear
-FEDformer
-Stationary
-Autoformer
-[12]
-[20]
-[5]
-[16]
-[15]
-[19]
-[18]
-[21]
-[11]
-[17]
-Metric
 MSE
@@
 96
-0.343 0.367 0.329 0.367
+0.343
+0.367 0.334 0.368
+0.355
+0.376
+0.329 0.367
 0.404
@@
 0.381
-0.383 0.367 0.385
+0.383 0.377
+0.391
+0.391
+0.392
+0.367 0.385
 0.450
@@
 0.414
-0.404 0.399 0.410
+0.404
+0.426
+0.420
+0.424
+0.415
+0.399 0.410
 0.532
@@
 720
-0.475 0.438 0.454 0.439
+0.475 0.438
+0.491
+0.459
+0.487
+0.450 0.454 0.439
 0.666
@@
 Avg
-0.403 0.398 0.387 0.400
+0.403 0.398
+0.407
+0.410
+0.414
+0.407
+0.387 0.400
 0.513
@@
 96
-0.181 0.264 0.175 0.259
+0.181 0.264 0.180 0.264
+0.182
+0.265 0.175 0.259
 0.287
@@
 0.305
-0.187 0.267
+0.187
+0.267
 0.189
@@
 192
-0.246 0.304 0.241 0.302
+0.246 0.304
+0.250
+0.309 0.246 0.304 0.241 0.302
 0.414
@@
 336
-0.307 0.343 0.305 0.343
+0.307 0.343 0.311 0.348 0.307 0.342 0.305 0.343
 0.597
@@
 0.321
-0.351 0.314
+0.351
+0.314
 0.357
@@
 720
+0.407 0.398
+0.412
+0.407
 0.407 0.398 0.402 0.400
@@
 Avg
-0.286 0.327 0.281 0.326
+0.286 0.327 0.288 0.332 0.286 0.327 0.281 0.326
 0.757
@@
 0.404
-0.291 0.333
+0.291
+0.333
 0.293
@@
 0.383 0.391
+0.386
+0.405
+0.386
+0.395
 0.414
@@
 0.464
-0.384 0.402
+0.384
+0.402
 0.494
@@
 0.433 0.419
+0.441
+0.436
+0.437
+0.424
 0.460
@@
 0.437
-0.432 0.420
+0.432
+0.420
 0.448
@@
 0.475 0.441
+0.487
+0.458 0.479 0.446
 0.501
@@
 0.488
-0.481 0.459 0.459 0.465
+0.481
+0.459
+0.459
+0.465
 0.588
@@
 720
-0.475 0.464 0.500 0.488
+0.475 0.464
+0.503
+0.491
+0.481 0.470 0.500 0.488
 0.653
@@
 0.441 0.429
+0.454
+0.447 0.446 0.434
 0.469
@@
 0.479
-0.456 0.452 0.440
+0.456
+0.452
+0.440
 0.460
@@
 96
-0.288 0.337 0.302 0.348
+0.288 0.337 0.297
+0.349
+0.288 0.338 0.302 0.348
 0.745
@@
 192
-0.375 0.392 0.388 0.400
+0.375 0.392 0.380 0.400 0.374 0.390
+0.388
+0.400
 0.877
@@
 0.509
-0.402 0.414
+0.402
+0.414
 0.430
@@
 336
-0.423 0.434 0.426 0.433
+0.423
+0.434
+0.428
+0.432 0.415 0.426 0.426 0.433
 1.043
@@
 0.571
-0.452 0.452
+0.452
+0.452
 0.485
@@
 720
-0.434 0.446 0.431 0.446
+0.434
+0.446 0.427 0.445 0.420 0.440 0.431 0.446
 1.104
@@
 0.679
-0.462 0.468
+0.462
+0.468
 0.500
@@
 Avg
-0.380 0.402 0.387 0.407
+0.380 0.402 0.383 0.407 0.374 0.398
+0.387
+0.407
 0.942
@@
 0.550
-0.414 0.427
+0.414
+0.427
 0.439
@@
 0.197
-0.273 0.181 0.270
+0.273
+0.148 0.240
+0.201
+0.281
+0.181
+0.270
 0.219
@@
 0.197
-0.276 0.188 0.274
+0.276 0.162 0.253
+0.201
+0.283
+0.188
+0.274
 0.231
@@
 0.212
-0.292 0.204 0.293
+0.292 0.178 0.269
+0.215
+0.298
+0.204
+0.293
 0.246
@@
 0.344
-0.198 0.300
+0.198
+0.300
 0.212
@@
 0.325
+0.225 0.317
+0.257
+0.331
 0.246
@@
 0.373
-0.220 0.320 0.233
+0.220 0.320
+0.233
 0.345
@@
 0.215
-0.292 0.205 0.290
+0.292 0.178 0.270
+0.219
+0.298
+0.205
+0.290
 0.244
@@
 0.344
-0.192 0.295
+0.192
+0.295
 0.208
@@
 96
-0.082 0.198
-0.177
+0.082 0.198 0.086
+0.206
+0.093
+0.217
+0.088
+0.205
+0.256
+0.367
+0.094
 0.218
-0.158
-0.230
-0.202
-0.261
 0.107
@@
 192
-0.176 0.297
-0.225
-0.259
-0.206
-0.277
-0.242
-0.298
+0.176 0.297 0.177 0.299
+0.184
+0.307
+0.176 0.299
+0.470
+0.509
+0.184
+0.307
 0.226
 0.344
-0.182
-0.303
+0.182 0.303
 0.215
@@
 336
-0.328
-0.412
-0.278 0.297 0.272 0.335 0.287 0.335
+0.328 0.412
+0.331
+0.417
+0.351
+0.432
+0.301 0.397
+1.268
+0.883
+0.349
+0.431
 0.367
@@
 0.700
-0.354 0.348 0.398 0.418 0.351 0.386
+0.847 0.691
+0.886
+0.714
+0.901
+0.714
+1.767
+1.068
+0.852
+0.698
 0.964
@@
 0.699
-0.839
-0.695
+0.839 0.695
 1.195
@@
 Avg
-0.362
-0.402
-0.259 0.281 0.259 0.315 0.271 0.320
+0.362 0.402 0.360 0.403
+0.378
+0.417
+0.367
+0.404
+0.940
+0.707
+0.370
+0.413
 0.416
@@
 0.383
-0.088 0.205 0.256
-0.367
-0.094 0.218
+0.395 0.268
+0.649
+0.389
+0.462 0.295 0.522 0.290
+0.805
+0.493
 0.593
@@
 0.360
-0.176 0.299 0.470
-0.509
-0.184 0.307
+0.417 0.276
+0.601
+0.366
+0.466 0.296 0.530 0.293
+0.756
+0.474
 0.617
@@
 0.362
-0.301
-0.397
-1.268
-0.883
-0.349
-0.431
+0.433 0.283
+0.609
+0.369
+0.482 0.304 0.558 0.305
+0.762
+0.477
 0.629
@@
 0.382
-0.901
-0.714
-1.767
-1.068
-0.852
-0.698
-0.640 0.350 0.632
+0.467 0.302
+0.647
+0.387
+0.514 0.322 0.589 0.328
+0.719
+0.449
+0.640
+0.350
+0.632
 0.396
@@
 0.623
-0.372 0.367
-0.404
-0.940
-0.707
-0.370
-0.413
+0.372
+0.428 0.282
+0.626
+0.378
+0.481 0.304 0.550 0.304
+0.760
+0.473
 0.620
@@
 96
-0.182 0.227
-0.462
-0.295
-0.522
-0.290
-0.805
-0.493
+0.182
+0.227
+0.174
+0.214
+0.192
+0.232
+0.177
+0.218 0.158
+0.230
+0.202
+0.261
 0.172 0.220
@@
 0.296
-0.173 0.223
+0.173
+0.223
 0.266
@@
 192
-0.227 0.262
-0.466
-0.296
-0.530
-0.293
-0.756
-0.474
+0.227
+0.262
+0.221 0.254
+0.240
+0.271
+0.225
+0.259 0.206
+0.277
+0.242
+0.298
 0.219 0.261
@@
 336
-0.279 0.298
-0.482
-0.304
-0.558
-0.305
-0.762
-0.477
+0.279 0.298 0.278 0.296
+0.292
+0.307
+0.278 0.297 0.272
+0.335
+0.287
+0.335
 0.280
@@
 720
-0.352
-0.344
-0.514
-0.322
-0.589
-0.328
-0.719
-0.449
+0.352 0.344
+0.358
+0.347
+0.364
+0.353
+0.354
+0.348
+0.398
+0.418
+0.351
+0.386
 0.365
 0.359
-0.352 0.288
-0.352
+0.352 0.288 0.352
 0.386
@@
 Avg
-0.260 0.283
-0.481
-0.304
-0.550
-0.304
-0.760
-0.473
-0.259 0.287
+0.260 0.283 0.258 0.278
+0.272
+0.291
+0.259 0.281 0.259
+0.315
+0.271
+0.320
+0.259
+0.287
 0.271
@@
 $1^{\text{st}}$ Count
-8 20 16 13
-1
+4
+15 10 15 5 6 12 5
+3
 0
-1
 0
-7 5
 0
@@
 1
+1
 0
+3
+0
+4
+0
+0
+0
+0
+0
+$2^{\text{nd}}$ Count 11 9 7
+15
 5
+8 7 17
+3
+0
+0
+0
+9
+0
 0
@@
 0
+3
 0
-$2^{\text{nd}}$ Count 13 7 7 15
-1
-3
-4
-3
-6 5
-2
-1
-0
-0
-2
-2
-0
-0
-5
 4
@@
 0
+0
+0
+0
 $3^{\text{rd}}$ Count
-5 5
-4
+8 5 9 9
+5
+6
+12
+11
 3
-3
-3
-2
-2
-11 20
-4
+0
+0
+0
+6 11
+0
+1
+1
 0
@@
 0
-6 4
-2
-1
+4
 0
-1
 0
-1
-B.4 Full Ablation Study
-We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting.
-We tested the forecasting length $L\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 6.
-In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases.
-There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored.
-According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 11.
-$l_{C+T}\leqslant min(l_{C},l_{T})$ (11)
-B.5 Full Hyperparameter Sensitivity
-The GPU memory in the experimental environment is limited, so we conducted a 2-dimensional traversal of the sample space size $|S|\in\{0,1,2,3,4,5,6\}$ and learning rate $lr\in\{1*10^{-2},5*10^{-2},1*10^{-1}\}$ in Table 8.
-In the case of sufficient GPU memory, more hyperparameters can be traversed.
-Based on the complete results of hyperparameter sensitivity, it can be seen that TEFN is an insensitive model to hyperparameters, which means that random hyperparameters can achieve stable and superior prediction performance.
-We use variance to describe the fluctuation of hyperparameters in Table 3.
-Table 8: Full results of hyperparameter sensitivity
+0
+0
+0
+Table 3. Full results of hyperparameter sensitivity
 $|S|$ $lr$ Target Dataset
@@
 720 0.255 0.326 0.479 0.467 0.473 0.479 0.507 0.457 0.427 0.414 1.055 0.783 0.645 0.384 0.376 0.368
+4.2. Comparative Baseline
+To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models.
+These models represent various approaches and techniques commonly used in time series forecasting, providing a robust benchmark for assessing TEFN’s capabilities.
+The selected baseline models are as follows:
+  • Transformer-based Models:
+    • iTransformer (ICLR 2024) (Liu et al., 2023): A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
+    • PatchTST (ICLR 2023) (Nie et al., 2023): This model divides the time series into patches and applies transformer-based operations on each patch independently.
+It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
+    • Crossformer (ICLR 2023) (Zhang and Yan, 2023): This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
+    • FEDformer (ICML 2022) (Zhou et al., 2022): A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
+    • ETSformer (Woo et al., 2022): A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
+    • Stationary (NeurIPS 2022) (Liu et al., 2022): A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
+    • Autoformer (NeurIPS 2021) (Wu et al., 2021): A self-supervised transformer model that learns to forecast future values without relying on explicit labels.
+It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
+  • Linear Models:
+    • Rlinear (ICLR 2024) (Li et al., 2023): This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data.
+It is effective at capturing periodic patterns and works well with longer input sequences.
+    • Dlinear (AAAI 2023) (Zeng et al., 2023): A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data.
+It serves as a simple yet effective baseline for comparison.
+    • LightTS (Zhang et al., 2022): A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
+  • Other Models:
+    • TiDE (ICLR 2023) (Das et al., 2023): A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
+    • TimeNet (ICLR 2023) (Wu et al., 2023): A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
+By comparing TEFN against this diverse set of baseline models, we can comprehensively assess its performance and identify its unique advantages and contributions to the field of time series forecasting.
+4.3. Implementation Details
+Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU.
+Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores , 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB.
+The proposed model TEFN is implemented in the PyTorch (Paszke et al., 2019).
+For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations.
+The open source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network.
+The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam (Kingma and Ba, 2015).
+We used Mean Absolute Error (MAE) and MSE in Equation 15 and Equation 16 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
+(15) $MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
+(16) $MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
+4.4. Forecasting Result
+In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table 2.
+The results showcase the remarkable capabilities of the Time Evidence Fusion Network (TEFN) in long-term time series forecasting.
+TEFN consistently achieves state-of-the-art (SOTA) performance across various datasets and prediction horizons, highlighting its effectiveness and versatility in capturing complex temporal patterns.
+Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer (Liu et al., 2023) and PatchTST (Nie et al., 2023), demonstrating its efficiency and scalability.
+TEFN has significant advantages over other models in terms of prediction accuracy, as evidenced by the MSE and MAE metrics.
+Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer (Liu et al., 2023) and $0.025$ smaller than PatchTST (Nie et al., 2023) in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset (Li et al., 2021).
+Similarly, in terms of MAE, TEFN once again showed a clear lead over iTransformer.
+TEFN was $0.008$ smaller than iTransformer (Liu et al., 2023) and $0.007$ smaller than PatchTST (Nie et al., 2023) in predicting a length of $L_{pred}=96$ on the Exchange dataset (Wu et al., 2021).
+These results highlight the effectiveness and robustness of TEFN in time series prediction, demonstrating its ability to provide more accurate predictions compared to contemporary methods.
+(a) TEFN (Ours)
+(b) PatchTST
+(c) Crossformer
+(d) TimesNet
+(e) Dlinear
+(f) Stationary
+Figure 5. Visualization of Electricity-96 tasks
+We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models.
+We take the task of Electricity-96 (ecl, [n.d.]) as an example and selected typical models PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), TimesNet (Wu et al., 2023), Dlinear (Zeng et al., 2023) and Stationary (Liu et al., 2022) for visualization in Figure 5.
+From the visualization results, TEFN fits the details and trends of time series better.
+4.5. Hyperparameter Sensitivity
+To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$.
+We take the Electricity-96 task as an example.
+TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 6.
+It is worth noting that the order of magnitude of the MSE in Figure 6 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small.
+Full hyperparameter sensitivity experiment results are in the Table 3.
+(a) $lr$ $|S|$MSE
+(b) Violin diagram of $|S|$ and MSE
+Figure 6. Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
+Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table 4.
+The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation.
+Random parameter selection within the traversal range of TEFN yields stable prediction results.
+Table 4. The variance of model error for traversal learning rate and sample space
+Metric Dataset
+($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
+MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
+192 0.0001 0.0000 0.0003 0.0004 0.0000 0.0000 0.0000 0.0002
+336 0.0001 0.0000 0.0002 0.0004 0.0000 0.0023 0.0000 0.0001
+720 0.0000 0.0000 0.0004 0.0001 0.0000 0.0038 0.0000 0.0001
+MAE 96 0.0000 0.0000 0.0001 0.0001 0.0001 0.0000 0.0000 0.0002
+192 0.0001 0.0000 0.0001 0.0001 0.0000 0.0000 0.0000 0.0002
+336 0.0001 0.0000 0.0001 0.0001 0.0000 0.0009 0.0000 0.0001
+720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
+(a) Channel BPA
+(b) Time BPA
+Figure 7.
+Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time.
+The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
+4.6. Ablation Study
+Table 5. The results of the ablation experiment
+Metric Target Model Dataset
+Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
+MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
+C 0.196$\downarrow$ 0.386$\uparrow$ 0.288 0.339$\downarrow$ 0.185$\uparrow$ 0.083$\uparrow$ 0.644$\downarrow$ 0.185$\uparrow$
+T 0.197 0.386$\uparrow$ 0.291$\uparrow$ 0.356$\uparrow$ 0.182$\uparrow$ 0.083$\uparrow$ 0.645 0.195$\uparrow$
+192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
+C 0.196$\downarrow$ 0.435$\uparrow$ 0.375 0.382$\uparrow$ 0.246 0.179$\uparrow$ 0.597$\downarrow$ 0.230$\uparrow$
+T 0.197 0.435$\uparrow$ 0.375 0.390$\uparrow$ 0.247$\uparrow$ 0.175$\downarrow$ 0.598 0.240$\uparrow$
+336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
+C 0.210$\downarrow$ 0.479$\uparrow$ 0.423 0.411$\downarrow$ 0.307 0.336$\uparrow$ 0.604$\downarrow$ 0.280$\uparrow$
+T 0.212 0.479$\uparrow$ 0.414$\downarrow$ 0.423$\uparrow$ 0.307 0.324$\downarrow$ 0.606$\uparrow$ 0.291$\uparrow$
+720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
+C 0.253 0.478$\uparrow$ 0.457$\uparrow$ 0.472$\downarrow$ 0.407 0.947$\uparrow$ 0.642$\downarrow$ 0.354$\uparrow$
+T 0.254$\uparrow$ 0.479$\uparrow$ 0.442$\uparrow$ 0.487$\uparrow$ 0.408$\uparrow$ 0.871$\uparrow$ 0.643$\downarrow$ 0.364$\uparrow$
+Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
+C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
+T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
+MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
+C 0.273$\downarrow$ 0.394$\uparrow$ 0.339$\uparrow$ 0.366$\downarrow$ 0.269$\uparrow$ 0.199$\uparrow$ 0.384 0.228$\uparrow$
+T 0.274 0.393$\uparrow$ 0.339$\uparrow$ 0.377$\uparrow$ 0.265$\uparrow$ 0.199$\uparrow$ 0.383$\downarrow$ 0.234$\uparrow$
+192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
+C 0.275$\downarrow$ 0.421$\uparrow$ 0.393$\uparrow$ 0.385$\uparrow$ 0.305$\uparrow$ 0.300$\uparrow$ 0.361$\uparrow$ 0.264$\uparrow$
+T 0.276 0.421$\uparrow$ 0.390$\downarrow$ 0.392$\uparrow$ 0.305$\uparrow$ 0.296$\downarrow$ 0.361$\uparrow$ 0.270$\uparrow$
+336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
+C 0.291$\downarrow$ 0.443$\uparrow$ 0.434$\downarrow$ 0.403$\downarrow$ 0.343 0.419$\uparrow$ 0.363 0.298
+T 0.292 0.445$\uparrow$ 0.425$\downarrow$ 0.413$\uparrow$ 0.342$\downarrow$ 0.410$\downarrow$ 0.362$\downarrow$ 0.306$\uparrow$
+720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
+C 0.325 0.463$\downarrow$ 0.462$\uparrow$ 0.436$\downarrow$ 0.399 0.739$\uparrow$ 0.382 0.347$\uparrow$
+T 0.324$\downarrow$ 0.469$\uparrow$ 0.454$\uparrow$ 0.449$\uparrow$ 0.398$\downarrow$ 0.702$\uparrow$ 0.382 0.354$\uparrow$
+Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
+C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
+T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
+$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
+TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$.
+In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table 5.
+Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error.
+Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (error of Electricty only varies $0.001$).
+It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
+We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting.
+We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 5.
+In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases.
+There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored.
+According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 17.
+(17) $l_{C+T}\leqslant min(l_{C},l_{T})$
+4.7. Efficiency Comparison
+TEFN is a compact model compared to the parameters of the Transformer.
+In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison.
+In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample.
+Different models occupy varying amounts of GPU and memory.
+Therefore, we used the size of the saved binary model files as the model size.
+We visualized this with a bubble chart in Figure 8 .
+It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters.
+Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency.
+TEFN maintains almost Dlinear (Zeng et al., 2023) performance and almost PatchTST (Nie et al., 2023) accuracy, making it a performance error balanced model.
+Figure 8.
+Model efficiency comparison.
+Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
+It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices.
+As evidenced by the official code repository (https://github.com/ztxtech/Time-Evidence-Fusion-Network), TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM.
+This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer (Zhang and Yan, 2023) when deployed on portable computers.
+4.8. Interpretability Analysis
+The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable.
+TEFN allocates mass to events within the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, where each event represents a potential outcome or scenario.
+The mass assigned to an event reflects the level of confidence or support for that event, providing insight into the model’s belief about its likelihood of occurrence.
+This support allocation is particularly interpretable due to the use of fuzzy membership functions $\mu(x)=wx+b$ within the BPA module, as depicted in Figure 7.
+By analyzing the shape and parameters of these membership functions, users can gain a clear understanding of the model’s perception of data features and their influence on predictions.
+For instance, it becomes evident which features are deemed more important than others, and how the model responds to variations within these features.
+This visualization not only makes the model’s inner workings more transparent but also allows users to intuitively assess the model’s confidence in its predictions and the level of uncertainty associated with each outcome.
+In summary, TEFN’s interpretability stems from its foundational reliance on fuzzy logic and the Dempster-Shafer theory, coupled with the transparency of its BPA and fuzzy reasoning processes.
+By exploring these aspects, users can develop a comprehensive understanding of the model’s decision-making process and prediction outcomes, fostering trust and confidence in the model’s capabilities.
+4.9. Limitation and Future Prospects
+TEFN is a naive and unmodified model.
+We lack skills in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method.
+Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well.
+In addition, for the fusion of mass distribution in the face of the failure of the classic Dempster Shafer rule for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern.
+In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels.
+In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution.
+Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic.
+Effective parameter initialization may exist, for example, initializing corresponding parameters in a non neural network way may lead to faster convergence.
+5. Conclusion
+This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets.
+TEFN treats different channels and time points as distinct information sources, using evidence theory to construct basic probability assignments and discard irrelevant data.
+This selective information fusion significantly reduces the computational burden, enabling TEFN to efficiently process and predict within acceptable time frames.
+The model’s parameter efficiency, robustness, and interpretability make it a valuable tool for time series forecasting tasks involving large and complex datasets.
+Acknowledgements.
+The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078)

### v3

@@
 Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting
-Abstract.
-In practical scenarios, time series forecasting necessitates timeliness, especially when dealing with large datasets.
+Abstract
+In practical scenarios, time series forecasting necessitates not only accuracy but also efficiency.
 Consequently, the exploration of model architectures remains a perennially trending topic in research.
-To meet these performance demands, we propose a novel backbone from the perspective of information fusion.
-Introducing the Basic Probability Assignment (BPA) Module and the Time Evidence Fusion Network (TEFN), based on evidence theory, allows us to achieve superior performance.
-On the other hand, the perspective of multi-source information fusion effectively improves the accuracy of forecasting.
-Due to the fact that BPA is generated by fuzzy theory, TEFN also has considerable interpretability.
-In real data experiments, the TEFN partially achieved state-of-the-art, with low errors comparable to PatchTST, and operating efficiency surpass performance models such as Dlinear.
-Meanwhile, TEFN has high robustness and small error fluctuations in the random hyperparameter selection.
-TEFN is not a model that achieves the ultimate in single aspect, but a model that balances performance, accuracy, stability, and interpretability.
-PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/ztxtech/Time-Evidence-Fusion-Network.
-1. Introduction
-The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains.
-From health diagnostics, where the progression of diseases is tracked over time, to financial analysis, where market trends are monitored for investment strategies, and weather forecasting, where atmospheric changes are predicted to safeguard communities, the importance of time series forecasting cannot be overstated.
-However, as the volume of data grows exponentially, time series forecasting faces significant performance challenges.
-The sheer scale of big data necessitates more efficient and faster algorithms to process and predict within acceptable time frames, which has become a pressing issue in the field.
-Figure 1.
-The channel dimension and time dimension of multivariate time series can be considered as different sources of information.
+To address these challenges, we propose a novel backbone architecture named Time Evidence Fusion Network (TEFN) from the perspective of information fusion.
+Specifically, we introduce the Basic Probability Assignment (BPA) Module based on evidence theory to capture the uncertainty of multivariate time series data from both channel and time dimensions.
+Additionally, we develop a novel multi-source information fusion method to effectively integrate the two distinct dimensions from BPA output, leading to improved forecasting accuracy.
+Lastly, we conduct extensive experiments to demonstrate that TEFN achieves performance comparable to state-of-the-art methods while maintaining significantly lower complexity and reduced training time.
+Also, our experiments show that TEFN exhibits high robustness, with minimal error fluctuations during hyperparameter selection.
+Furthermore, due to the fact that BPA is derived from fuzzy theory, TEFN offers a high degree of interpretability.
+Therefore, the proposed TEFN balances accuracy, efficiency, stability, and interpretability, making it a desirable solution for time series forecasting.
+Index Terms:
+Time Series, Forecasting, Evidence Theory, Information Fusion, Neural Network
+publicationid: pubid: $0$–$00$ 2021 IEEE
+1 Introduction
+The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1], such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11].
+The importance of time series forecasting cannot be overstated.
+There are also many methods for time series forecasting.
+The initial methods came from statistics, such as Autoregressive (AR) model, Moving Average (MA) model, Autoregressive Moving Average (ARMA) model, and Autoregressive integrated moving average (ARIMA) model [12].
+Deep learning is a powerful method with wide applications in various disciplines [13, 14].
+In recent years, time series prediction models based on neural networks have become a research hotspot.
+These neural networks have evolved from the simplest deep neural networks (DNN) [15].
+Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on.
+Large Language Models (LLM) for time series also perform well in cross dataset forecasting tasks [22, 23, 24].
+There is uncertainty in time series data, and modeling uncertainty on time series is effective.
+Evidence theory is an uncertain reasoning framework based on mass function [25, 26, 27, 28].
+Previous studies have shown that evidence theory can effectively utilize the uncertainty of time-series data and achieve positive application results [29, 30].
+On the other hand, evidence theory is also a method of information fusion that can effectively aggregate different information sources and make effective decisions [31, 32].
+Meanwhile, evidence theory can also be applied in neural networks to accomplish tasks such as classification [33].
+Therefore, in order to reduce the error of time series forecasting, we propose the Time Evidence Fusion Network (TEFN) from the perspectives of evidence theory and information fusion.
+TEFN is a lightweight neural network that considers the time and channel dimensions of multivariate time series as different sources of information in Figure 1.
+Figure 1: The channel dimension and time dimension of multivariate time series can be considered as different sources of information.
 Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
-In response to these challenges, researchers have been focusing on optimizing the structure of the backbone models used in time series forecasting.
-It has been observed that linear-based models, such as Dlinear (Zeng et al., 2023), often exhibit superior performance compared to transformer-based models like Crossformer (Zhang and Yan, 2023), particularly in terms of computational efficiency and speed.
-These findings highlight the potential of structural optimizations in enhancing the capabilities of time series forecasting models.
-Nevertheless, the quest for an ideal model that balances performance, prediction accuracy, stability, and interpretability continues.
-The aim is to develop a model that can provide robust support for high-stakes scenarios where timely and reliable predictions are critical, thereby bridging the gap between theoretical advancements and practical application needs.
-Evidence theory, also known as Dempster-Shafer theory, provides a mathematical framework for representing and combining degrees of belief in the presence of uncertainty (Shafer, 1976; DEMPSTER, 1967).
-It is particularly useful in situations where information is incomplete, ambiguous, or comes from diverse sources.
-The theory is grounded in the concept of a belief function, which quantifies the degree to which a hypothesis is supported by the available evidence, without necessarily committing to a probability distribution over the possible outcomes.
-One of the key strengths of evidence theory is its ability to handle the fusion of multiple information sources effectively.
-This makes it a powerful tool in fields such as data fusion, sensor networks, and multi-criteria decision-making, where diverse pieces of information need to be integrated to form a coherent union.
-In the realm of symbolic logic, evidence theory often appears as a reasoning framework that offers high interpretability.
-Unlike traditional probability theory, which requires precise probability assignments, evidence theory allows for a more nuanced representation of uncertainty.
-It does this through the use of belief and plausibility measures, which provide a range of values for the likelihood of an event occurring, rather than a single probability.
-An innovative application of evidence theory in machine learning is the development of evidence neural networks by Denoeux (Denoeux, 2000).
-These networks are designed for classification tasks and integrate the principles of evidence theory into the architecture of neural networks.
-By doing so, they enhance the network’s ability to handle uncertainty and provide a more robust and interpretable decision-making process.
-The evidence neural networks output belief masses that can be directly interpreted as measures of evidence supporting each class, thus offering a clear understanding of the network’s reasoning process.
-Therefore, we introduce the Time Evidence Fusion Network (TEFN), a novel architecture designed specifically for regression tasks within the realm of time series forecasting.
-The conceptual underpinning of TEFN is the innovative strategy of dissecting the time series data into dual perspectives—time and channels—thereby treating each channel and temporal instance as a distinct and autonomous source of information.
-This approach allows us to leverage the principles of evidence theory to fuse these diverse information streams within an integrated framework, aiming to enhance the accuracy and reliability of predictions as depicted in Figure 1.
-In essence, TEFN is built upon the premise of simplicity and naivety in its design philosophy.
-It refrains from incorporating complex technical modifications that often render models opaque and difficult to interpret.
-Instead, TEFN maintains a straightforward structure that prioritizes clarity and ease of understanding, making it a transparent and accessible backbone for time series analysis.
-By adopting this minimalist approach, TEFN emerges as a refreshing alternative in the field, offering a new foundation upon which more sophisticated forecasting models can potentially be built, while retaining the interpretability and trustworthiness that are hallmarks of evidence theory.
-In the architecture of TEFN, we introduce a pivotal component termed the Basic Probability Assignment (BPA) Module, which stands as the core of our evidence-theoretic approach to time series forecasting.
-The BPA Module operates on the fundamental principles of evidence theory, meticulously mapping various sources of information into mass distributions that are intrinsically linked to the desired target outcome.
-This mapping is achieved through the application of fuzzy membership functions, which assign membership grades to different fuzzy sets, thereby constructing a comprehensive target distribution as proposed by Deng (Deng, 2015).
-In contrast to the convolutional operations prevalent in image processing, which selectively condense features pertinent to the target through filters (Bengio et al., 1993), the BPA Module embodies an expansionary process, as illustrated in Figure 2.
-Rather than narrowing down to specific features, BPA embraces an expansive perspective, contemplating all potentialities within the event space, which is a collection of subsets derived from the sample space.
-This expansion across various events is particularly advantageous for the simplistic structures inherent in time series data, enabling a thorough exploration of the myriad hidden informational layers within.
-For the diverse distributions yielded by the multitude of information sources, we treat each source as an individual sample, and by amalgamating these samples, we utilize the expected forecasting values as the ultimate prediction, thereby delivering a robust and informed forecast that encapsulates the collective wisdom of all contributing sources.
-Figure 2. Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
-In this paper, we delve into the practical nuances of the Time Evidence Fusion Network (TEFN), examining its precision, memory consumption, training velocity, sensitivity to parameters, and interpretability.
-To maintain equitable comparisons, all experimental data and configuration files are sourced from open-source initiatives.
-TEFN has partially achieved state-of-the-art (SOTA) performance on various real-world datasets, with a notable reduction in training duration and model complexity when contrasted with Transformer-based models such as Crossformer (Zhang and Yan, 2023) and Patch-based models like PatchTST (Nie et al., 2023).
-In terms of random hyperparameter selection, TEFN demonstrates minimal variability, affirming its stability.
-The contributions of this manuscript are summarized as follows:
-  • The proposed method is well-suited for large datasets, showcasing exceptional performance.
+In TEFN, we propose a novel Basic Probability Assignment (BPA) Module.
+Based on the evidence theory of decision-making process, BPA maps different sources of information to distributions related to the target outcome.
+The target distribution is achieved by the fuzzy membership of different fuzzy sets [34].
+In image processing, convolution condenses the required features of the target [35].
+Unlike convolution, BPA is an expansion process in Figure 2.
+On the contrary, BPA considers all possibilities in the event space composed of subsets of the sample space, expanding on different events, which is more effective for simple structures like time series and can fully explore the different hidden information inside.
+For different distributions generated by multiple information sources, we consider sources as samples and use the expected forecasting values as the final result.
+Figure 2: Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
+In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability.
+At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects.
+TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20, 36].
+For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model.
+The contributions of this article are as follows:
   • A novel evidential backbone BPA designed for extracting potential distributions from simple data structures is introduced.
@@
 Finally, we conclude the paper by summarizing our key findings in the Section 5.
-2. Related Work
-2.1. Dempster-Shafer Theory
-Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory (Shafer, 1976; DEMPSTER, 1967).
+2 Related Work
+2.1 Dempster-Shafer Theory
+Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [25, 26].
 It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities.
-When combining two mass functions, denoted as $m_{1}$ and $m_{2}$, the resulting fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation 1.
-The symbols $A$, $B$, and $C$ represent focal elements, which are typically combinations of all independent samples (or target labels, such as the classification targets).
-(1) $m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$
-2.2. Basic Probability Assignment
+All independent and mutually exclusive sample sets that exist in the system are called frames of discrete (FOD), represented as $\Omega$.
+When there are multiple sources of information, information fusion is usually required to make decisions.
+Fusing two mass functions, denoted as $m_{1}$ and $m_{2}$, the fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation (1).
+The symbol variables $A$, $B$, and $C$ are focal elements, which are typically combinations of all independent samples in the FOD, which are often composed of target labels, such as the classification targets.
+$m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
+2.2 Basic Probability Assignment
 BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions.
 Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making.
-The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation 2 where $w$ and $b$ is weight and bias of the membership function.
-BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation 3.
-(2) $\begin{split}m=\mu(x)=wx+b\end{split}$
-(3) $\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$
+The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation (2) where $w$ and $b$ is weight and bias of the membership function.
+BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (3).
+$\begin{split}m=\mu(x)=wx+b\end{split}$ (2)
+$\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$ (3)
 The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information.
-2.3. Evidence Decision Making
-The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) (Deng, 2015; Zhan et al., 2024).
+2.3 Evidence Decision Making
+The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [34, 37].
 BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution.
 Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution.
-The most classic fusion method is the DSR (Shafer, 1976; DEMPSTER, 1967) mentioned in Section 2.1.
+The most classic fusion method is the DSR [25, 26] mentioned in Section 2.1.
 Finally, decision making transpires as the progression from the attained distribution towards the designated target.
-The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignostic probability transformation (PPT) (Dubois et al., 1993) in Equation 4, where $x$ is a target label and $e$ are combinations of all possible target labels.
-(4) $p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$
-Figure 3. The overall structure of TEFN
-2.4. Linear Time Series Model
-There are many linear time series forecasting models (LTSFM).
-LTSFM offer several advantages for modeling and forecasting time series data.
-Firstly, their simplicity and transparency make them easy to implement, interpret, and analyze.
-The parameters of linear models directly represent the relationships between input features and the output, providing insights into the underlying dynamics of the time series.
-Secondly, LTSFM exhibit high efficiency, requiring fewer computational resources and hyperparameters compared to complex models like transformer-based models (Vaswani, 2017; Zhang and Yan, 2023).
+The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignistic probability transformation (PPT) [38] in Equation (4), where $x$ is a target label and $e$ are combinations of all possible target labels.
+$p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$ (4)
+2.4 Linear Time Series Model
+There are many linear time series forecasting models (LTSFM) [12, 39, 40, 41]. LTSFM offer several advantages for modeling and forecasting time series data:
+  • The simplicity and transparency make them easy to implement, interpret, and analyze.
   • This leads to faster training and inference times, making them suitable for real-time applications and large datasets.
-A classic LTSFM model is Dlinear (Zeng et al., 2023).
+  • No additional encoding layers are required.
+A classic LTSFM model is Dlinear [39].
 DLinear serves as a simple yet effective baseline for long-term time series forecasting.
@@
   • TEFN is a general framework that does not require identification and prediction of trends and periodicity, while Dlinear is more like a feature modeling prediction of time series.
-3. Time Evidence Fusion Network
+3 Time Evidence Fusion Network
 The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels.
@@
 The overall structure of TEFN is depicted in Figure 3.
-Figure 4.
-The operation process of a BPA module for a time series $T$ of length $L=5$.
-The sample space size in the BPA module is set as $|S|=2$, and the event space size is set as $|2^{S}|=2^{|S|}=4$.
-The operation of the BPA module is similar to that of the convolution module, but the BPA module diffuses different masses instead of summing them in the convolution.
-3.1. Time Normalization and De-normalization
-To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model (Liu et al., 2022).
-This process computes the mean $\mu$ as shown in Equation 5 and variance $\sigma^{2}$ as shown in Equation 6 for each segment of the time series $x$.
-The subsequent normalization and de-normalization operations, presented in Equation 7 and Equation 8, respectively, act as symmetrical transformations that standardize the time series data.
+Next, we will elaborate on each module one by one.
+Figure 3: The overall structure of TEFN: Establish corresponding mass functions for the time dimension and channel dimension as the target dimensions of BPA.
+Figure 4: The operation of a BPA module: weight matrix and bias matrix are both learnable parameters. BPA elevates the dimensionality of data by adding a dimension to distinguish different events.
+3.1 Time Normalization and De-normalization
+To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model [42].
+This process computes the mean $\mu$ as shown in Equation (5) and variance $\sigma^{2}$ as shown in Equation (6) for each segment of the time series $x$.
+The subsequent normalization and de-normalization operations, presented in Equation (7) and Equation (8), where $\hat{y}$ is forecasting result, respectively, act as symmetrical transformations that standardize the time series data.
 The neural network operation $Net(\cdot)$ is then applied to the normalized time series $x_{Norm}$, ensuring that the model learns from data with consistent scale and distribution.
 This normalization strategy effectively mitigates the influence of outliers and initial data scale, fostering faster convergence and improved stability during training.
-(5) $\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$
-(6) $\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$
-(7) $x_{Norm}=\frac{x-\mu}{\sigma}$
-(8) $x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$
-3.2. Time Dimension Projection
-In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation 9.
+$\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$ (5)
+$\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$ (6)
+$x_{Norm}=\frac{x-\mu}{\sigma}$ (7)
+$\hat{y}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (8)
+3.2 Time Dimension Projection
+In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation (9).
 Here, $L_{pred}$ represents the desired length of the prediction horizon.
-This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet (Wu et al., 2023).
-(9) $x^{\prime}=project(x)=x\times W_{p}+b$
+This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet [43].
+$x^{\prime}=project(x)=x\times W_{p}+b$ (9)
 By extending the input vector to match the prediction length, the time projection layer facilitates the direct generation of future values based on the learned temporal patterns.
 This approach simplifies the model design and enables efficient prediction without the need for iterative or recursive computations, as commonly observed in autoregressive models.
-3.3. Basic Probability Assignment Module
+3.3 Basic Probability Assignment Module
 In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data.
@@
 This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data.
-For classification problems, $S$ corresponds to all the possible labels, while for regression problems, $S$ is treated as a set of anonymous samples, ensuring mutual exclusivity among them.
-This representation allows us to capture the fuzzy nature of the data, where each dimension generates a mass distribution within the event space.
-To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation 10.
+Since time series forecasting is a regression problem, the size of the sample space needs to be used as a hyperparameter.
+This representation allows us to capture the fuzzy features of the input time series, where each dimension generates a mass distribution within the event space.
+To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation (10).
 Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$.
 The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$.
-The fuzzy membership function $\mu_{k}$ is defined in Equation 10 where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function, respectively.
+The fuzzy membership function $\mu_{k}$ is defined in Equation (10) where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function.
 These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space.
-(10) $m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$
+Figure 4 shows the visible process of BPA module.
+$m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (10)
 By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions.
@@
 To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module.
-Instead of the linear form presented in Equation 10, we can adopt non-linear functions that capture more complex relationships within the data.
+Instead of the linear form presented in Equation (10), we can adopt non-linear functions that capture more complex relationships within the data.
 Several options exist for non-linear membership functions, including:
-  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation 3, capturing the likelihood of each data point belonging to different membership levels.
+  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation (3), capturing the likelihood of each data point belonging to different membership levels.
 The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
-  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation 11.
+  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation (11).
 The sigmoid function can capture both gradual and abrupt changes in membership values.
-    (11) $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$
-  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation 12 where $g(x)$ is a non-linear function.
+    $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$ (11)
+  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation (12) where $g(x)$ is a non-linear function.
 This approach is useful for capturing complex patterns and relationships within the data.
-    (12) $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$
+    $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$ (12)
 By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data.
 This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
-3.4. Expectation Fusion
+3.4 Expectation Fusion
 TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$.
 This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$.
-To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation 13.
+To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation (13).
 Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
-(13) $\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$
-The fusion operation in Equation 13 can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions.
+$\begin{split}\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})&=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}\\ &=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}\end{split}$ (13)
+The fusion operation in Equation (13) can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions.
 The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
-TEFN deliberately avoids using the Dempster-Shafer Rule of Combination (DSR) for several reasons.
-Firstly, the computational complexity of DSR is significantly higher compared to the expectation-based fusion method.
+TEFN deliberately avoids using the DSR for several reasons:
+  1. 1.
+    The computational complexity of DSR is significantly higher compared to the expectation-based fusion method.
 Additionally, DSR involves point-to-point multiplication of mass distributions within the same dimension, which can lead to accuracy loss when dealing with high-dimensional data due to finite precision limitations.
-Secondly, DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1.
+  2. 2.
+    DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1.
 This sensitivity can result in the final fused result being dominated by that single distribution, potentially leading to conflicts and loss of information from other distributions.
@@
 This approach effectively integrates information from different dimensions while mitigating the risks of conflicts and accuracy loss associated with DSR.
-The final normalized prediction results are then aggregated across dimensions using the equation in Equation 14 and restored to their original scale through the de-normalization layer.
-(14) $\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$
-4. Experiments
+The final normalized prediction results are then aggregated across dimensions using the equation in Equation (14) and restored to their original scale through the de-normalization layer.
+$\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (14)
+4 Experiments
 The experimental section aims to comprehensively evaluate TEFN’s performance in long-term time series forecasting.
@@
 Lastly, we will explore TEFN’s interpretability by analyzing the basic probability assignments generated by its BPA module.
-4.1. Dataset
+4.1 Dataset
 Our research employs the TEFN for the challenging task of long-term time series prediction.
@@
 The selected datasets are:
-  • Electricity (ecl, [n.d.]): This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time.
+  • Electricity [44]: This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time.
 It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
-  • ETT (4 subsets) (Li et al., 2021): This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges.
+  • ETT (4 subsets) [19]: This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges.
 It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
-  • Exchange (Wu et al., 2021): This dataset captures exchange rate fluctuations between major currencies.
+  • Exchange [45]: This dataset captures exchange rate fluctuations between major currencies.
 It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
-  • Traffic (tra, [n.d.]): This dataset contains traffic flow data collected from sensors in a city.
+  • Traffic [46]: This dataset contains traffic flow data collected from sensors in a city.
 It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
-  • Weather (wea, [n.d.]): This dataset includes historical weather data, encompassing various meteorological parameters.
+  • Weather [47]: This dataset includes historical weather data, encompassing various meteorological parameters.
 It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
@@
 The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability.
-More dataset features are in the Table 1.
-Table 1. The quantitative features of the dataset
+More dataset features are in the Table I.
+TABLE I: The quantitative features of the dataset
 Dataset Dimension Dataset Size
@@
 Exchange 8 (5120, 665, 1422)
-Table 2.
-This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks.
+TABLE II: This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks.
 The models are evaluated on four different prediction lengths: $L_{pred}\in\{96,192,336,720\}$, providing a thorough assessment of their performance under varying forecasting horizons.
@@
 Autoformer
+[48]
+[49]
+[36]
+[20]
+[50]
+[43]
+[51]
+[40]
+[39]
+[52]
+[42]
+[45]
 Metric
@@
 0.381
-0.383 0.377
+0.383
+0.377
 0.391
@@
 0.404
-0.374 0.387
+0.374
+0.387
 0.408
@@
 0.425
-0.410 0.411
+0.410
+0.411
 0.435
@@
 720
-0.475 0.438
+0.475
+0.438
 0.491
@@
 0.487
-0.450 0.454 0.439
+0.450
+0.454 0.439
 0.666
@@
 Avg
-0.403 0.398
+0.403
+0.398
 0.407
@@
 0.419
-0.400 0.406
+0.400
+0.406
 0.429
@@
 96
-0.181 0.264 0.180 0.264
+0.181
+0.264 0.180 0.264
 0.182
-0.265 0.175 0.259
+0.265
+0.175 0.259
 0.287
@@
 0.250
-0.309 0.246 0.304 0.241 0.302
+0.309
+0.246 0.304 0.241 0.302
 0.414
@@
 0.364
-0.249 0.309
+0.249
+0.309
 0.253
@@
 336
-0.307 0.343 0.311 0.348 0.307 0.342 0.305 0.343
+0.307 0.343
+0.311
+0.348
+0.307 0.342 0.305 0.343
 0.597
@@
 0.524
-0.408 0.403
+0.408
+0.403
 0.414
@@
 Avg
-0.286 0.327 0.288 0.332 0.286 0.327 0.281 0.326
+0.286 0.327
+0.288
+0.332
+0.286 0.327 0.281 0.326
 0.757
@@
 0.386
-0.400 0.376
+0.400
+0.376
 0.419
@@
 0.492
-0.436 0.429
+0.436
+0.429
 0.538
@@
 0.487
-0.458 0.479 0.446
+0.458
+0.479
+0.446
 0.501
@@
 0.491
-0.481 0.470 0.500 0.488
+0.481 0.470
+0.500
+0.488
 0.653
@@
 0.454
-0.447 0.446 0.434
+0.447
+0.446
+0.434
 0.469
@@
 0.349
-0.288 0.338 0.302 0.348
+0.288 0.338
+0.302
+0.348
 0.745
@@
 192
-0.375 0.392 0.380 0.400 0.374 0.390
+0.375 0.392
+0.380
+0.400
+0.374 0.390
 0.388
@@
 0.428
-0.432 0.415 0.426 0.426 0.433
+0.432 0.415 0.426
+0.426
+0.433
 1.043
@@
 0.434
-0.446 0.427 0.445 0.420 0.440 0.431 0.446
+0.446
+0.427 0.445 0.420 0.440
+0.431
+0.446
 1.104
@@
 Avg
-0.380 0.402 0.383 0.407 0.374 0.398
+0.380 0.402
+0.383
+0.407
+0.374 0.398
 0.387
@@
 0.329
-0.168 0.272
+0.168
+0.272
 0.187
@@
 0.197
-0.276 0.162 0.253
+0.276
+0.162 0.253
 0.201
@@
 0.325
-0.225 0.317
+0.225
+0.317
 0.257
@@
 0.355
-0.222 0.321
+0.222
+0.321
 0.254
@@
 0.215
-0.292 0.178 0.270
+0.292
+0.178 0.270
 0.219
@@
 96
-0.082 0.198 0.086
+0.082 0.198
+0.086
 0.206
@@
 0.344
-0.182 0.303
+0.182
+0.303
 0.215
@@
 336
-0.328 0.412
+0.328
+0.412
 0.331
@@
 0.700
-0.847 0.691
+0.847
+0.691
 0.886
@@
 Avg
-0.362 0.402 0.360 0.403
+0.362
+0.402 0.360 0.403
 0.378
@@
 0.389
-0.462 0.295 0.522 0.290
+0.462
+0.295
+0.522
+0.290
 0.805
@@
 0.366
-0.466 0.296 0.530 0.293
+0.466
+0.296
+0.530
+0.293
 0.756
@@
 0.369
-0.482 0.304 0.558 0.305
+0.482 0.304
+0.558
+0.305
 0.762
@@
 0.387
-0.514 0.322 0.589 0.328
+0.514 0.322
+0.589
+0.328
 0.719
@@
 0.378
-0.481 0.304 0.550 0.304
+0.481 0.304
+0.550
+0.304
 0.760
@@
 0.261
-0.172 0.220
+0.172
+0.220
 0.197
@@
 0.262
-0.221 0.254
+0.221
+0.254
 0.240
@@
 0.298
-0.219 0.261
+0.219
+0.261
 0.237
@@
 336
-0.279 0.298 0.278 0.296
+0.279
+0.298
+0.278 0.296
 0.292
@@
 720
-0.352 0.344
+0.352
+0.344
 0.358
@@
 0.359
-0.352 0.288 0.352
+0.352
+0.288
+0.352
 0.386
@@
 Avg
-0.260 0.283 0.258 0.278
+0.260
+0.283
+0.258 0.278
 0.272
@@
 4
-15 10 15 5 6 12 5
+15 10 15
+5
+6 12
+5
 3
@@
 0
-$2^{\text{nd}}$ Count 11 9 7
-15
+$2^{\text{nd}}$ Count 11 9
+7
+6
 5
-8 7 17
+8
+7
+17
+1
 3
+1
+0
+9
+1
+1
+1
+0
+0
+3
+1
+0
+0
+2
 0
@@
 0
-9
-0
-0
-1
-1
-0
-3
-0
-4
-0
-0
-0
-0
-0
-$3^{\text{rd}}$ Count
-8 5 9 9
-5
-6
-12
-11
-3
-0
-0
-0
-6 11
-0
-1
-1
-0
-3
-0
-4
-0
-0
-0
-0
-0
-Table 3. Full results of hyperparameter sensitivity
-$|S|$ $lr$ Target Dataset
-Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-0 $1*10^{-2}$ 96 0.197 0.273 0.403 0.407 0.288 0.337 0.344 0.367 0.182 0.264 0.086 0.204 0.645 0.384 0.186 0.230
-192 0.197 0.276 0.433 0.420 0.375 0.392 0.385 0.385 0.247 0.305 0.178 0.299 0.598 0.360 0.232 0.266
-336 0.212 0.292 0.478 0.443 0.423 0.434 0.414 0.404 0.310 0.344 0.328 0.412 0.605 0.362 0.283 0.301
-720 0.254 0.326 0.476 0.465 0.452 0.458 0.478 0.438 0.410 0.400 0.886 0.711 0.644 0.382 0.357 0.348
-$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.289 0.338 0.344 0.368 0.182 0.265 0.082 0.198 0.645 0.384 0.182 0.227
-192 0.197 0.277 0.433 0.420 0.376 0.393 0.383 0.385 0.247 0.305 0.179 0.300 0.598 0.360 0.232 0.266
-336 0.213 0.294 0.476 0.441 0.425 0.437 0.415 0.406 0.308 0.344 0.340 0.421 0.605 0.363 0.284 0.301
-720 0.254 0.327 0.478 0.465 0.453 0.459 0.478 0.439 0.408 0.399 0.911 0.722 0.644 0.382 0.352 0.344
-$1*10^{-1}$ 96 0.197 0.274 0.383 0.391 0.291 0.340 0.347 0.369 0.182 0.264 0.082 0.198 0.645 0.383 0.183 0.228
-192 0.197 0.277 0.433 0.419 0.376 0.394 0.383 0.385 0.257 0.314 0.180 0.300 0.598 0.360 0.244 0.274
-336 0.214 0.295 0.476 0.441 0.424 0.435 0.416 0.406 0.309 0.344 0.340 0.421 0.605 0.363 0.292 0.308
-720 0.255 0.327 0.482 0.469 0.454 0.460 0.477 0.439 0.408 0.400 0.937 0.732 0.644 0.382 0.352 0.345
-1 $1*10^{-2}$ 96 0.197 0.274 0.384 0.392 0.289 0.337 0.349 0.370 0.182 0.265 0.083 0.199 0.645 0.385 0.186 0.229
-192 0.197 0.277 0.438 0.421 0.377 0.392 0.382 0.386 0.246 0.304 0.178 0.299 0.598 0.361 0.232 0.265
-336 0.212 0.292 0.475 0.441 0.425 0.435 0.418 0.407 0.307 0.343 0.335 0.418 0.606 0.363 0.280 0.299
-720 0.253 0.325 0.476 0.464 0.452 0.458 0.482 0.442 0.409 0.400 0.883 0.709 0.644 0.382 0.355 0.347
-$5*10^{-2}$ 96 0.198 0.275 0.386 0.393 0.294 0.341 0.346 0.369 0.182 0.265 0.088 0.205 0.645 0.385 0.194 0.239
-192 0.197 0.277 0.438 0.422 0.375 0.392 0.383 0.385 0.247 0.305 0.184 0.304 0.598 0.360 0.241 0.274
-336 0.212 0.292 0.475 0.441 0.453 0.448 0.414 0.404 0.308 0.343 0.361 0.434 0.605 0.363 0.282 0.300
-720 0.255 0.327 0.475 0.465 0.458 0.461 0.478 0.439 0.407 0.398 0.886 0.708 0.644 0.383 0.352 0.345
-$1*10^{-1}$ 96 0.198 0.275 0.388 0.395 0.295 0.342 0.345 0.369 0.182 0.265 0.089 0.206 0.645 0.383 0.202 0.249
-192 0.197 0.277 0.434 0.420 0.376 0.394 0.381 0.384 0.247 0.305 0.176 0.297 0.598 0.360 0.239 0.276
-336 0.212 0.292 0.476 0.441 0.423 0.435 0.415 0.404 0.309 0.343 0.343 0.423 0.605 0.363 0.289 0.306
-720 0.256 0.328 0.476 0.466 0.463 0.465 0.478 0.439 0.408 0.398 0.892 0.710 0.644 0.383 0.360 0.353
-2 $1*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.290 0.338 0.347 0.369 0.183 0.266 0.083 0.200 0.645 0.383 0.187 0.230
-192 0.197 0.276 0.434 0.420 0.375 0.393 0.386 0.386 0.247 0.305 0.177 0.298 0.598 0.360 0.232 0.266
-336 0.212 0.292 0.476 0.441 0.423 0.435 0.414 0.405 0.311 0.346 0.348 0.427 0.605 0.362 0.284 0.301
-720 0.253 0.325 0.476 0.466 0.453 0.459 0.478 0.439 0.408 0.399 0.898 0.719 0.644 0.382 0.355 0.347
-$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.289 0.338 0.345 0.368 0.182 0.265 0.084 0.202 0.646 0.384 0.183 0.228
-192 0.197 0.277 0.434 0.420 0.377 0.394 0.384 0.387 0.247 0.305 0.177 0.298 0.598 0.360 0.228 0.263
-336 0.212 0.292 0.477 0.443 0.424 0.436 0.415 0.405 0.308 0.343 0.353 0.430 0.605 0.363 0.293 0.309
-720 0.254 0.325 0.478 0.466 0.458 0.461 0.477 0.439 0.408 0.400 0.930 0.732 0.644 0.382 0.355 0.347
-$1*10^{-1}$ 96 0.197 0.274 0.384 0.391 0.289 0.339 0.347 0.368 0.182 0.265 0.088 0.205 0.646 0.385 0.186 0.231
-192 0.197 0.277 0.438 0.423 0.378 0.395 0.387 0.387 0.247 0.305 0.191 0.310 0.598 0.361 0.241 0.278
-336 0.212 0.293 0.480 0.446 0.425 0.437 0.416 0.406 0.308 0.343 0.403 0.463 0.605 0.363 0.299 0.316
-720 0.254 0.326 0.475 0.466 0.463 0.465 0.478 0.439 0.410 0.402 0.880 0.707 0.644 0.382 0.357 0.350
-3 $1*10^{-2}$ 96 0.197 0.274 0.385 0.393 0.290 0.339 0.343 0.368 0.182 0.265 0.083 0.199 0.645 0.383 0.183 0.227
-192 0.197 0.277 0.434 0.420 0.376 0.393 0.383 0.385 0.248 0.306 0.183 0.304 0.598 0.360 0.232 0.266
-336 0.212 0.293 0.477 0.443 0.426 0.435 0.414 0.405 0.308 0.344 0.338 0.420 0.606 0.363 0.282 0.300
-720 0.253 0.325 0.477 0.465 0.475 0.473 0.476 0.438 0.407 0.399 0.943 0.739 0.644 0.382 0.353 0.346
-$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.290 0.337 0.344 0.368 0.183 0.266 0.083 0.200 0.645 0.383 0.189 0.236
-192 0.198 0.277 0.435 0.421 0.382 0.398 0.384 0.385 0.258 0.316 0.200 0.317 0.598 0.360 0.234 0.269
-336 0.213 0.293 0.476 0.441 0.426 0.437 0.415 0.406 0.308 0.344 0.353 0.429 0.606 0.363 0.285 0.303
-720 0.254 0.325 0.489 0.474 0.460 0.463 0.478 0.439 0.412 0.402 1.031 0.775 0.644 0.383 0.357 0.350
-$1*10^{-1}$ 96 0.197 0.274 0.384 0.391 0.291 0.339 0.346 0.369 0.184 0.267 0.084 0.201 0.645 0.384 0.196 0.245
-192 0.198 0.277 0.437 0.423 0.381 0.398 0.385 0.386 0.248 0.306 0.176 0.297 0.598 0.360 0.254 0.290
-336 0.212 0.292 0.475 0.441 0.430 0.440 0.416 0.406 0.328 0.361 0.362 0.435 0.607 0.363 0.288 0.308
-720 0.254 0.325 0.492 0.477 0.434 0.446 0.478 0.440 0.418 0.405 1.046 0.782 0.644 0.383 0.357 0.350
-4 $1*10^{-2}$ 96 0.197 0.274 0.393 0.399 0.288 0.338 0.344 0.368 0.185 0.268 0.084 0.201 0.645 0.383 0.182 0.227
-192 0.197 0.276 0.434 0.420 0.376 0.393 0.381 0.385 0.247 0.305 0.179 0.300 0.598 0.360 0.227 0.262
-336 0.212 0.292 0.478 0.444 0.424 0.436 0.415 0.405 0.307 0.343 0.338 0.420 0.606 0.363 0.279 0.298
-720 0.255 0.326 0.476 0.465 0.457 0.462 0.477 0.439 0.409 0.399 0.889 0.713 0.644 0.383 0.353 0.345
-$5*10^{-2}$ 96 0.197 0.274 0.383 0.392 0.292 0.340 0.345 0.368 0.181 0.264 0.085 0.203 0.645 0.384 0.203 0.254
-192 0.198 0.278 0.433 0.420 0.377 0.394 0.385 0.386 0.248 0.307 0.179 0.300 0.598 0.360 0.234 0.270
-336 0.213 0.293 0.486 0.449 0.425 0.435 0.417 0.407 0.312 0.347 0.346 0.427 0.607 0.363 0.294 0.314
-720 0.257 0.330 0.480 0.467 0.462 0.464 0.479 0.440 0.408 0.399 0.861 0.700 0.644 0.383 0.359 0.353
-$1*10^{-1}$ 96 0.198 0.274 0.384 0.392 0.293 0.342 0.358 0.376 0.183 0.266 0.085 0.203 0.646 0.384 0.211 0.262
-192 0.197 0.277 0.434 0.421 0.397 0.411 0.399 0.395 0.249 0.307 0.180 0.300 0.598 0.360 0.255 0.292
-336 0.215 0.295 0.489 0.452 0.473 0.467 0.420 0.407 0.309 0.344 0.377 0.447 0.606 0.363 0.305 0.324
-720 0.255 0.328 0.485 0.470 0.454 0.460 0.480 0.441 0.414 0.404 0.923 0.727 0.644 0.383 0.375 0.364
-5 $1*10^{-2}$ 96 0.198 0.275 0.384 0.392 0.291 0.341 0.343 0.367 0.182 0.265 0.083 0.201 0.646 0.385 0.184 0.229
-192 0.198 0.277 0.434 0.421 0.379 0.396 0.381 0.385 0.247 0.305 0.187 0.306 0.600 0.362 0.227 0.262
-336 0.212 0.293 0.476 0.442 0.434 0.443 0.416 0.406 0.308 0.343 0.340 0.421 0.606 0.364 0.280 0.299
-720 0.254 0.326 0.479 0.466 0.457 0.462 0.479 0.440 0.409 0.401 0.957 0.744 0.645 0.384 0.354 0.348
-$5*10^{-2}$ 96 0.198 0.274 0.383 0.391 0.300 0.348 0.349 0.372 0.183 0.267 0.085 0.204 0.645 0.386 0.204 0.249
-192 0.198 0.278 0.433 0.419 0.378 0.395 0.382 0.383 0.248 0.306 0.177 0.299 0.599 0.362 0.251 0.288
-336 0.212 0.292 0.475 0.441 0.428 0.439 0.418 0.407 0.313 0.348 0.363 0.437 0.607 0.365 0.301 0.320
-720 0.265 0.339 0.478 0.467 0.504 0.488 0.478 0.441 0.409 0.400 0.986 0.755 0.645 0.384 0.357 0.351
-$1*10^{-1}$ 96 0.198 0.275 0.386 0.393 0.316 0.360 0.375 0.388 0.197 0.280 0.084 0.201 0.647 0.388 0.209 0.260
-192 0.199 0.279 0.434 0.420 0.418 0.425 0.393 0.390 0.251 0.310 0.178 0.299 0.600 0.362 0.280 0.307
-336 0.241 0.330 0.476 0.442 0.454 0.457 0.435 0.417 0.326 0.359 0.541 0.541 0.606 0.363 0.293 0.314
-720 0.278 0.355 0.479 0.469 0.505 0.494 0.505 0.456 0.421 0.410 1.052 0.780 0.645 0.385 0.375 0.366
-6 $1*10^{-2}$ 96 0.199 0.276 0.384 0.392 0.290 0.340 0.347 0.369 0.183 0.265 0.095 0.214 0.649 0.389 0.194 0.240
-192 0.199 0.279 0.437 0.422 0.379 0.396 0.383 0.385 0.247 0.305 0.191 0.311 0.601 0.365 0.230 0.266
-336 0.213 0.294 0.478 0.444 0.430 0.440 0.417 0.406 0.308 0.343 0.372 0.443 0.609 0.367 0.285 0.303
-720 0.257 0.329 0.481 0.466 0.469 0.468 0.475 0.438 0.411 0.402 0.915 0.724 0.648 0.387 0.355 0.348
-$5*10^{-2}$ 96 0.198 0.274 0.383 0.391 0.298 0.345 0.379 0.387 0.185 0.268 0.094 0.215 0.648 0.386 0.204 0.250
-192 0.198 0.277 0.436 0.421 0.416 0.422 0.388 0.386 0.248 0.307 0.181 0.303 0.599 0.361 0.252 0.289
-336 0.213 0.294 0.475 0.442 0.433 0.439 0.426 0.412 0.310 0.346 0.439 0.487 0.606 0.363 0.288 0.308
-720 0.254 0.326 0.487 0.473 0.512 0.491 0.482 0.441 0.414 0.405 0.935 0.733 0.645 0.384 0.360 0.355
-$1*10^{-1}$ 96 0.198 0.275 0.384 0.392 0.341 0.380 0.396 0.401 0.223 0.305 0.088 0.206 0.647 0.386 0.214 0.263
-192 0.234 0.320 0.439 0.423 0.427 0.430 0.478 0.441 0.260 0.319 0.182 0.304 0.599 0.361 0.253 0.289
-336 0.258 0.335 0.481 0.444 0.435 0.442 0.507 0.459 0.325 0.359 0.372 0.445 0.607 0.364 0.307 0.325
-720 0.255 0.326 0.479 0.467 0.473 0.479 0.507 0.457 0.427 0.414 1.055 0.783 0.645 0.384 0.376 0.368
-4.2. Comparative Baseline
+4.2 Comparative Baseline
 To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models.
@@
   • Transformer-based Models:
-    • iTransformer (ICLR 2024) (Liu et al., 2023): A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
-    • PatchTST (ICLR 2023) (Nie et al., 2023): This model divides the time series into patches and applies transformer-based operations on each patch independently.
+    • iTransformer (ICLR 2024) [48]: A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
+    • PatchTST (ICLR 2023) [36]: This model divides the time series into patches and applies transformer-based operations on each patch independently.
 It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
-    • Crossformer (ICLR 2023) (Zhang and Yan, 2023): This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
-    • FEDformer (ICML 2022) (Zhou et al., 2022): A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
-    • ETSformer (Woo et al., 2022): A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
-    • Stationary (NeurIPS 2022) (Liu et al., 2022): A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
-    • Autoformer (NeurIPS 2021) (Wu et al., 2021): A self-supervised transformer model that learns to forecast future values without relying on explicit labels.
+    • Crossformer (ICLR 2023) [20]: This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
+    • FEDformer (ICML 2022) [52]: A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
+    • ETSformer [51]: A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
+    • Stationary (NeurIPS 2022) [42]: A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
+    • Autoformer (NeurIPS 2021) [45]: A self-supervised transformer model that learns to forecast future values without relying on explicit labels.
 It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
   • Linear Models:
-    • Rlinear (ICLR 2024) (Li et al., 2023): This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data.
+    • Rlinear (ICLR 2024) [49]: This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data.
 It is effective at capturing periodic patterns and works well with longer input sequences.
-    • Dlinear (AAAI 2023) (Zeng et al., 2023): A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data.
+    • Dlinear (AAAI 2023) [39]: A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data.
 It serves as a simple yet effective baseline for comparison.
-    • LightTS (Zhang et al., 2022): A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
+    • LightTS [40]: A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
   • Other Models:
-    • TiDE (ICLR 2023) (Das et al., 2023): A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
-    • TimeNet (ICLR 2023) (Wu et al., 2023): A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
+    • TiDE (ICLR 2023) [50]: A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
+    • TimesNet (ICLR 2023) [43]: A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
 By comparing TEFN against this diverse set of baseline models, we can comprehensively assess its performance and identify its unique advantages and contributions to the field of time series forecasting.
-4.3. Implementation Details
+4.3 Implementation Details
 Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU.
 Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores, 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB.
-The proposed model TEFN is implemented in the PyTorch (Paszke et al., 2019).
-For the sake of fairness, we used an open-source benchmark located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations.
-The open source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network.
-The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam (Kingma and Ba, 2015).
-We used Mean Absolute Error (MAE) and MSE in Equation 15 and Equation 16 as error indicators in the experiment where $y_{i}$ is actual value and $\hat{y}_{i}$ is predicted value.
-(15) $MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
-(16) $MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
-4.4. Forecasting Result
-In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table 2.
+The proposed model TEFN is implemented in the PyTorch [53].
+For the sake of fairness, we used an open-source benchmark  [^1] The open-source benchmark Time-Series-Library located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations.
+The open-source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network..
+The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam [54].
+We used Mean Absolute Error (MAE) and MSE in Equation (15) and Equation (16) as error indicators in the experiment where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the predicted value.
+$MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (15)
+$MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (16)
+4.4 Forecasting Result
+In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table II.
 The results showcase the remarkable capabilities of the Time Evidence Fusion Network (TEFN) in long-term time series forecasting.
 TEFN consistently achieves state-of-the-art (SOTA) performance across various datasets and prediction horizons, highlighting its effectiveness and versatility in capturing complex temporal patterns.
-Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer (Liu et al., 2023) and PatchTST (Nie et al., 2023), demonstrating its efficiency and scalability.
+Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer [48] and PatchTST [36], demonstrating its efficiency and scalability.
 TEFN has significant advantages over other models in terms of prediction accuracy, as evidenced by the MSE and MAE metrics.
-Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer (Liu et al., 2023) and $0.025$ smaller than PatchTST (Nie et al., 2023) in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset (Li et al., 2021).
+Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer [48] and $0.025$ smaller than PatchTST [36] in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset [19].
 Similarly, in terms of MAE, TEFN once again showed a clear lead over iTransformer.
-TEFN was $0.008$ smaller than iTransformer (Liu et al., 2023) and $0.007$ smaller than PatchTST (Nie et al., 2023) in predicting a length of $L_{pred}=96$ on the Exchange dataset (Wu et al., 2021).
+TEFN was $0.008$ smaller than iTransformer [48] and $0.007$ smaller than PatchTST [36] in predicting a length of $L_{pred}=96$ on the Exchange dataset [45].
 These results highlight the effectiveness and robustness of TEFN in time series prediction, demonstrating its ability to provide more accurate predictions compared to contemporary methods.
-(a) TEFN (Ours)
-(b) PatchTST
-(c) Crossformer
-(d) TimesNet
-(e) Dlinear
-(f) Stationary
-Figure 5. Visualization of Electricity-96 tasks
+(a) TEFN (Ours)-Electricity
+(b) PatchTST-Electricity
+(c) Crossformer-Electricity
+(d) TimesNet-Electricity
+(e) Dlinear-Electricity
+(f) Stationary-Electricity
+(g) TEFN (Ours)-ETTm1
+(h) PatchTST-ETTm1
+(i) Crossformer-ETTm1
+(j) TimesNet-ETTm1
+(k) Dlinear-ETTm1
+(l) Stationary-ETTm1
+Figure 5: Visualization of long-term forecasting tasks: Figure 5(a) - Figure 5(f) in Electricity dataset with $L_{pred}=96$ and Figure 5(g) - Figure 5(l) in ETTm1 dataset with $L_{pred}=96$
 We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models.
-We take the task of Electricity-96 (ecl, [n.d.]) as an example and selected typical models PatchTST (Nie et al., 2023), Crossformer (Zhang and Yan, 2023), TimesNet (Wu et al., 2023), Dlinear (Zeng et al., 2023) and Stationary (Liu et al., 2022) for visualization in Figure 5.
+We take the task of Electricity-96 [44] and ETTm1-96 as an example and selected typical models PatchTST [36], Crossformer [20], TimesNet [43], Dlinear [39] and Stationary [42] for visualization in Figure 5.
 From the visualization results, TEFN fits the details and trends of time series better.
-4.5. Hyperparameter Sensitivity
+4.5 Efficiency Comparison
+TEFN is a compact model compared to the parameters of the Transformer.
+In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison.
+In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample.
+Different models occupy varying amounts of GPU and memory.
+Therefore, we used the size of the saved binary model files as the model size.
+We visualized this with a bubble chart in Figure 6 .
+It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters.
+Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency.
+TEFN maintains almost Dlinear [39] performance and almost PatchTST [36] accuracy, making it a performance error balanced model.
+Figure 6: Model efficiency comparison.
+Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iteration.
+The models beyond the canvas have not been drawn.
+It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices.
+As evidenced by the official code repository  [^2] https://github.com/ztxtech/Time-Evidence-Fusion-Network, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM.
+This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer [20] when deployed on portable computers.
+According to the process of BPA, the time complexity of BPA is $O(n*2^{|S|})$ where $n$ is input size and $|S|$ is sample space size.
+For the TEFN model, the specific time complexity is $O(C*L*2^{|S|})$ where $C$ is channel size, and $L$ is the length of time series.
+$C$ and $|S|$ are constants and satisfy $C,|S|<<L$.
+Hence, the final time complexity is $O(L)$.
+So TEFN can maintain positive performance in long-term forecasting tasks.
+Another factor affecting experimental efficiency is whether TEFN is linear or nonlinear.
+We chose to use a linear BPA module in the experimental section because of its high efficiency.
+Nonlinear modules also exist, and only the membership functions need to be replaced.
+BPA needs to correspond to different types of nonlinear time series and design membership functions.
+In the future, we will explore adaptive nonlinear functions.
+The fitting of nonlinear functions also brings higher computational costs, therefore, in the case of efficiency priority, the linear version of TEFN is the preferred choice.
+4.6 Hyperparameter Sensitivity
+(a) $lr$ $|S|$MSE
+(b) Violin diagram of $|S|$ and MSE
+Figure 7: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
 To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$.
@@
 It is worth noting that the order of magnitude of the MSE in Figure 7 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small.
-Full hyperparameter sensitivity experiment results are in the Table 3.
-(a) $lr$ $|S|$MSE
-(b) Violin diagram of $|S|$ and MSE
-Figure 6. Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
-Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table 4.
+In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [55, 56]  [^3] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/).
+is conducted on prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 8.
+According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE.
+Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table III.
 The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation.
 Random parameter selection within the traversal range of TEFN yields stable prediction results.
-Table 4. The variance of model error for traversal learning rate and sample space
+TABLE III: The variance of model error for traversal learning rate and sample space
 Metric Dataset
@@
 720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
-(a) Channel BPA
-(b) Time BPA
-Figure 7.
-Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time.
-The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
-4.6. Ablation Study
-Table 5. The results of the ablation experiment
+Figure 8: Heatmap of hyperparameter sensitivity correlation
+4.7 Ablation Study
+TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$.
+In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table IV.
+Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error.
+Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (the error of Electricity only varies $0.001$).
+It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
+We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting.
+We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table IV.
+In the aspect of all the results of ablation, TEFN can effectively improve performance by combining the dimensions of time $T$ and channel $C$ in most cases.
+There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored.
+According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
+$l_{C+T}\leqslant min(l_{C},l_{T})$ (17)
+TABLE IV: The results of the ablation experiment
 Metric Target Model Dataset
@@
 $\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
-TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$.
-In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table 5.
-Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error.
-Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (error of Electricty only varies $0.001$).
-It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
-We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting.
-We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table 5.
-In the aspect of all the results of ablation, TEFN can effectively improve the performance by combining the dimensions of time $T$ and channel $C$ in most cases.
-There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored.
-According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation 17.
-(17) $l_{C+T}\leqslant min(l_{C},l_{T})$
-4.7. Efficiency Comparison
-TEFN is a compact model compared to the parameters of the Transformer.
-In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison.
-In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample.
-Different models occupy varying amounts of GPU and memory.
-Therefore, we used the size of the saved binary model files as the model size.
-We visualized this with a bubble chart in Figure 8 .
-It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters.
-Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency.
-TEFN maintains almost Dlinear (Zeng et al., 2023) performance and almost PatchTST (Nie et al., 2023) accuracy, making it a performance error balanced model.
-Figure 8.
-Model efficiency comparison.
-Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iterations.
-It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices.
-As evidenced by the official code repository (https://github.com/ztxtech/Time-Evidence-Fusion-Network), TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM.
-This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer (Zhang and Yan, 2023) when deployed on portable computers.
-4.8. Interpretability Analysis
+(a) Channel BPA
+(b) Time BPA
+Figure 9: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time.
+The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
+4.8 Robustness Experiments
+TEFN is a highly robust neural network.
+To verify the robustness of TEFN in real prediction, a comparative experiment was conducted by adding noise to the dataset.
+The unit noise $\gamma$ follows the Gaussian distribution $N(0,1)$ with an expectation of 00 and a variance of $1$.
+Assume that the input time series is $x$, with time and channel dimensions.
+To adapt to the scale of the data, the transformation as Eq.(18) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function.
+So, the noisy data experiments are compared with the original data experiments.
+$x^{\prime}_{ij}=x_{ij}+std(x_{i})*\gamma$ (18)
+The experimental results of robustness testing are shown in the Table V.
+Continuing from the analysis, the experimental outcomes underscore the robust nature of TEFN, where the introduction of Gaussian noise to the dataset led to only marginal increases in prediction errors, with most variations being insignificant relative to the original errors.
+Notably, in long-term forecasting scenarios, a counterintuitive improvement in accuracy was observed in some instances, indicating that TEFN’s adaptability and learning capacity in complex, dynamic environments extend beyond mere resilience to noise, potentially leveraging subtle perturbations to enhance prediction performance.
+TABLE V: The results of robustness experiments
+Dataset Pred Original Noisy
+MSE MAE MSE MAE
+ETTm1 96 0.343 0.367 0.414 ($\uparrow$ 0.071) 0.411 ($\uparrow$ 0.044)
+192 0.381 0.383 0.443 ($\uparrow$ 0.062) 0.429 ($\uparrow$ 0.046)
+336 0.414 0.404 0.499 ($\uparrow$ 0.085) 0.458 ($\uparrow$ 0.054)
+720 0.475 0.438 0.381 ($\downarrow$ 0.094) 0.398 ($\downarrow$ 0.040)
+ETTm2 96 0.181 0.264 0.254 ($\uparrow$ 0.073) 0.312 ($\uparrow$ 0.048)
+192 0.246 0.304 0.314 ($\uparrow$ 0.068) 0.350 ($\uparrow$ 0.046)
+336 0.307 0.343 0.411 ($\uparrow$ 0.104) 0.403 ($\uparrow$ 0.060)
+720 0.407 0.398 0.199 ($\downarrow$ 0.208) 0.281 ($\downarrow$ 0.117)
+ETTh1 96 0.383 0.391 0.466 ($\uparrow$ 0.083) 0.448 ($\uparrow$ 0.057)
+192 0.433 0.419 0.508 ($\uparrow$ 0.075) 0.468 ($\uparrow$ 0.049)
+336 0.475 0.441 0.509 ($\uparrow$ 0.034) 0.484 ($\uparrow$ 0.043)
+720 0.475 0.464 0.426 ($\downarrow$ 0.049) 0.427 ($\downarrow$ 0.037)
+ETTh2 96 0.288 0.337 0.386 ($\uparrow$ 0.098) 0.401 ($\uparrow$ 0.064)
+192 0.375 0.392 0.432 ($\uparrow$ 0.057) 0.441 ($\uparrow$ 0.049)
+336 0.423 0.434 0.466 ($\uparrow$ 0.043) 0.467 ($\uparrow$ 0.033)
+720 0.434 0.446 0.304 ($\downarrow$ 0.130) 0.352 ($\downarrow$ 0.094)
+Electricity 96 0.197 0.273 0.270 ($\uparrow$ 0.073) 0.363 ($\uparrow$ 0.090)
+192 0.197 0.276 0.281 ($\uparrow$ 0.084) 0.371 ($\uparrow$ 0.095)
+336 0.212 0.292 0.322 ($\uparrow$ 0.110) 0.399 ($\uparrow$ 0.107)
+720 0.253 0.325 0.270 ($\uparrow$ 0.017) 0.361 ($\uparrow$ 0.036)
+Traffic 96 0.645 0.383 0.727 ($\uparrow$ 0.082) 0.479 ($\uparrow$ 0.096)
+192 0.598 0.360 0.734 ($\uparrow$ 0.136) 0.481 ($\uparrow$ 0.121)
+336 0.605 0.362 0.768 ($\uparrow$ 0.163) 0.491 ($\uparrow$ 0.129)
+720 0.644 0.382 0.774 ($\uparrow$ 0.130) 0.494 ($\uparrow$ 0.112)
+Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.050) 0.269 ($\uparrow$ 0.042)
+192 0.227 0.262 0.285 ($\uparrow$ 0.058) 0.305 ($\uparrow$ 0.043)
+336 0.279 0.298 0.357 ($\uparrow$ 0.078) 0.350 ($\uparrow$ 0.052)
+720 0.352 0.344 0.190 ($\downarrow$ 0.162) 0.237 ($\downarrow$ 0.107)
+Exchange 96 0.082 0.198 0.188 ($\uparrow$ 0.106) 0.308 ($\uparrow$ 0.110)
+192 0.176 0.297 0.340 ($\uparrow$ 0.164) 0.422 ($\uparrow$ 0.125)
+336 0.328 0.412 0.963 ($\uparrow$ 0.635) 0.739 ($\uparrow$ 0.327)
+720 0.861 0.700 0.092 ($\downarrow$ 0.769) 0.213 ($\downarrow$ 0.487)
+4.9 Interpretability Analysis
 The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable.
@@
 By exploring these aspects, users can develop a comprehensive understanding of the model’s decision-making process and prediction outcomes, fostering trust and confidence in the model’s capabilities.
-4.9. Limitation and Future Prospects
+4.10 Limitation and Future Prospects
 TEFN is a naive and unmodified model.
-We lack skills in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method.
+There is not engineering experience in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method.
 Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well.
-In addition, for the fusion of mass distribution in the face of the failure of the classic Dempster Shafer rule for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern.
+In addition, for the fusion of mass distribution in the face of the failure of the classic DSR for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern.
 In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels.
@@
 Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic.
-Effective parameter initialization may exist, for example, initializing corresponding parameters in a non neural network way may lead to faster convergence.
-5. Conclusion
+Effective parameter initialization may exist, for example, initializing corresponding parameters in a non-neural network way may lead to faster convergence.
+5 Conclusion
 This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets.
 TEFN treats different channels and time points as distinct information sources, using evidence theory to construct basic probability assignments and discard irrelevant data.
-This selective information fusion significantly reduces the computational burden, enabling TEFN to efficiently process and predict within acceptable time frames.
+This selective information fusion significantly reduces the computational burden, enabling TEFN to process and predict within acceptable time frames efficiently.
 The model’s parameter efficiency, robustness, and interpretability make it a valuable tool for time series forecasting tasks involving large and complex datasets.
-Acknowledgements.
-The work is partially supported by National Natural Science Foundation of China (Grant No. 62373078)

### v4

@@
 Time Series, Forecasting, Evidence Theory, Information Fusion, Neural Network
-publicationid: pubid: $0$–$00$ 2021 IEEE
-1 Introduction
-The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1], such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11].
+I Introduction
+The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1] such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11].
 The importance of time series forecasting cannot be overstated.
-There are also many methods for time series forecasting.
-The initial methods came from statistics, such as Autoregressive (AR) model, Moving Average (MA) model, Autoregressive Moving Average (ARMA) model, and Autoregressive integrated moving average (ARIMA) model [12].
+There are many methods for time series forecasting.
+The initial methods came from statistics, such as Autoregressive integrated moving average (ARIMA) model [12].
 Deep learning is a powerful method with wide applications in various disciplines [13, 14].
@@
 These neural networks have evolved from the simplest deep neural networks (DNN) [15].
-Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on.
+Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on [22, 23].
 Large Language Models (LLM) for time series also perform well in cross dataset forecasting tasks [24, 25, 26].
@@
 Evidence theory is an uncertain reasoning framework based on mass function [27, 28, 29, 30].
-Previous studies have shown that evidence theory can effectively utilize the uncertainty of time-series data and achieve positive application results [29, 30].
+Previous studies have shown that evidence theory can effectively utilize the uncertainty of time series data and achieve positive application results [31, 32].
 On the other hand, evidence theory is also a method of information fusion that can effectively aggregate different information sources and make effective decisions [33, 34].
 Meanwhile, evidence theory can also be applied in neural networks to accomplish tasks such as classification [35].
+Furthermore, the essence of evidence theory lies in using mass functions to model variables.
+A mass function, as a generalized form of probability, can directly represent the uncertainty among different arguments that is challenging to achieve with traditional probability models.
 Therefore, in order to reduce the error of time series forecasting, we propose the Time Evidence Fusion Network (TEFN) from the perspectives of evidence theory and information fusion.
@@
 For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model.
+TEFN holds significant value and potential impact in various practical application scenarios.
+In the energy sector, such as power systems, accurate electricity consumption forecasting is crucial for the reliable operation, management, and planning of the power grid [39, 40].
+With large scale power consumption data, traditional forecasting methods may face high computational complexity and long training times.
+TEFN, however, can efficiently process millions of time series, enabling power companies to optimize power generation and distribution, reducing energy waste and costs.
+In the context of AIOps, resource forecasting and optimization require accurate predictions to ensure the stable operation of cloud systems [41].
+As cloud platforms manage millions of time series data related to resource usage, low complexity and high efficiency forecasting methods are essential [42].
+TEFN can provide precise resource usage forecasts, helping cloud operators to allocate resources effectively, reduce operational costs, and improve service reliability.
 The contributions of this article are as follows:
@@
 This paper is organized as follows.
-We begin by reviewing relevant literature in the Section 2.
-Next, in the Section 3, we delve into the architecture and underlying mathematical principles of TEFN.
-In the Section 4, we present a series of experiments that demonstrate TEFN’s superior performance.
-Finally, we conclude the paper by summarizing our key findings in the Section 5.
-2 Related Work
-2.1 Dempster-Shafer Theory
+We begin by reviewing relevant literature in the Section II.
+Next, in the Section III, we delve into the architecture and underlying mathematical principles of TEFN.
+In the Section IV, we present a series of experiments that demonstrate TEFN’s superior performance.
+Finally, we conclude the paper by summarizing our key findings in the Section V.
+II Related Work
+II-A Dempster-Shafer Theory
 Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [27, 28].
 It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities.
-All independent and mutually exclusive sample sets that exist in the system are called frames of discrete (FOD), represented as $\Omega$.
+All independent and mutually exclusive sample sets that exist in the system are called frames of discernment (FOD), represented as $\Omega$.
 When there are multiple sources of information, information fusion is usually required to make decisions.
@@
 $m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
-2.2 Basic Probability Assignment
+Here is an example of an intuitive explanation of the mass function.
+Suppose there are two unidentifiable features $A$ and $B$ in the time series.
+There are different ways $1$ and $2$ to identify whether it is $A$ or $B$ (For example, whether the increase in time series values is due to a trend or a cycle.).
+Therefore, we obtain the results in in Equation (2), and both ways are able to distinguish the feature of $A$, but there is also a certain degree that couldn’t distinguish between $A$ and $B$.
+It should be noted that if modeling is done using probabilities, the inability to distinguish between $A,B$ cannot be directly expressed.
+The usual approach is to declare a new argument $C$ as indistinguishable between $A$ and $B$.
+But at this time, $C$ will be independent of $A$ and $B$, and the connection between propositions will be lost.
+DSR can unify the opinions of the two methods $m_{1},m_{2}$ and obtain a comprehensive opinion $m$ in Equation (2).
+[…]
+II-B Basic Probability Assignment
 BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions.
@@
 The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation (3) where $w$ and $b$ is weight and bias of the membership function.
-BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (3).
+This kind of membership is called the triangular membership function, assuming that the membership is linearly related to the random variable.
 $\begin{split}m=\mu(x)=wx+b\end{split}$ (3)
-$\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$ (3)
+An intuitive example is shown in Figure 3.
+There are two fuzzy sets, which are assumed to represent the categories of random variables.
+We randomly take a sample $x_{i}$, and at this time, the membership function reflects the degree to which $x_{i}$ belongs to this fuzzy set, which can be understood as the degree of belonging to this category.
+This is also different from probability.
+The membership degree is bounded and can only reflect the extent to which it belongs to the current fuzzy set.
+Therefore, when multiple fuzzy sets appear, they need to be normalized in order to be converted into a probability distribution or a mass function.
+And the mass function can increase the description of uncertainty.
+In Figure 3, since the degree of belonging to the first category is greater, we believe that the degree of belonging to the second category actually reflects that the two fuzzy sets are not clearly distinguishable.
+Figure 3: The conversion between fuzzy membership and mass function.
+BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (taking probability density function as a membership function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (4).
+[…]
 The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information.
-2.3 Evidence Decision Making
+On the one hand, the existence of fuzzy sets allows elements to partially belong to a set, which allows for a more continuous attribution, rather than the set only belonging to this set or not belonging to this set.
+II-C Evidence Decision Making
 The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [36, 43].
@@
 Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution.
-The most classic fusion method is the DSR [25, 26] mentioned in Section 2.1.
+The most classic fusion method is the DSR [27, 28] mentioned in Section II-A.
 Finally, decision making transpires as the progression from the attained distribution towards the designated target.
 The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignistic probability transformation (PPT) [44] in Equation (5), where $x$ is a target label and $e$ are combinations of all possible target labels.
-$p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$ (4)
-2.4 Linear Time Series Model
+$p(x)=\sum_{x\in e}\frac{m(e)}{|e|}$ (5)
+II-D Linear Time Series Model
 There are many linear time series forecasting models (LTSFM) [12, 45, 46, 47]. LTSFM offer several advantages for modeling and forecasting time series data:
@@
 The final prediction is obtained by summing the outputs of the two linear regressions.
-TEFN shares similarities with Dlinear in that both models utilize linear regression for prediction. However, there are key differences:
-  • DLinear employs a simpler decomposition method compared to TEFN, which may limit its ability to capture intricate temporal dynamics.
-  • DLinear focuses on capturing linear relationships and trends in the data, while TEFN focuses on the mining of internal hidden relationships, so it strengthens predictions through dimensionality increase and uncertain reasoning.
-  • TEFN is a general framework that does not require identification and prediction of trends and periodicity, while Dlinear is more like a feature modeling prediction of time series.
-3 Time Evidence Fusion Network
+III Time Evidence Fusion Network
 The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels.
@@
 Figure 5: The operation of a BPA module: weight matrix and bias matrix are both learnable parameters. BPA elevates the dimensionality of data by adding a dimension to distinguish different events.
-3.1 Time Normalization and De-normalization
+III-A Time Normalization and De-normalization
 To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model [48].
@@
 $\hat{y}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (9)
-3.2 Time Dimension Projection
+III-B Time Dimension Projection
 In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation (10).
@@
 This approach simplifies the model design and enables efficient prediction without the need for iterative or recursive computations, as commonly observed in autoregressive models.
-3.3 Basic Probability Assignment Module
+III-C Basic Probability Assignment Module
 In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data.
@@
 Since time series forecasting is a regression problem, the size of the sample space needs to be used as a hyperparameter.
+It can be considered that the BPA operation can be regarded as an increase in dimension, and the size of the new dimension is the same as the power set.
+This operation expands the different representations of the original elements on the power set.
+The simplest form of the generating function is a linear function, which can be understood as a fuzzy membership function.
+Of course, it can also be a nonlinear function.
+In neural networks, a nonlinear layer is usually added after the linear transformation.
 This representation allows us to capture the fuzzy features of the input time series, where each dimension generates a mass distribution within the event space.
@@
 The fuzzy membership function $\mu_{k}$ is defined in Equation (11) where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function.
+The intuitive understanding of Equation (11) is that the different values $j$ on the input data dimension $D$ are expanded into power sets, where $k$ corresponds to different elements in the power set.
 These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space.
-Figure 4 shows the visible process of BPA module.
 $m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (11)
+Figure 5 shows the visible process of BPA module in time dimension.
+It should be noted that this is not matrix multiplication, but a multiplication of the same number of different parameters for each element, and the results are combined into a new dimension.
+The values in the Figure 5 corresponding to different time steps are extended to the mass of the fuzzy set memship function measure in the event space.
+These membership functions are shared among different channels.
+Therefore, the size of the newly added dimension is $2^{|S|}$, reflecting the mass function of different events.
+The operation on the channel dimension is symmetrical, and it shares parameters for different time steps.
 By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions.
 This approach provides a principled way to handle the inherent fuzziness and ambiguity in real-world time series, leading to improved performance and reliability in various applications.
+Unlike the Bayesian layer and the probability layer, BPA considers imprecise probabilities, allowing for the uncertainty of the probability modeling itself, rather than using probabilities as a basis to estimate uncertainty.
+An intuitive example is the DSR case in Section II-A.
 To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module.
 Instead of the linear form presented in Equation (11), we can adopt non-linear functions that capture more complex relationships within the data.
-Several options exist for non-linear membership functions, including:
-  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation (3), capturing the likelihood of each data point belonging to different membership levels.
-The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
-  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation (11).
-The sigmoid function can capture both gradual and abrupt changes in membership values.
-    $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$ (11)
-  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation (12) where $g(x)$ is a non-linear function.
-This approach is useful for capturing complex patterns and relationships within the data.
-    $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$ (12)
+Several options exist for non-linear membership functions, including ReLU Function, Tanh Function, Piecewise Functions, etc.
 By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data.
 This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
-3.4 Expectation Fusion
+III-D Expectation Fusion
 TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$.
@@
 Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
-$\begin{split}\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})&=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}\\ &=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}\end{split}$ (13)
 The fusion operation in Equation (12) can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions.
 The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
+$\begin{split}\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})&=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}\\ &=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}\end{split}$ (12)
 TEFN deliberately avoids using the DSR for several reasons:
@@
 $\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (13)
-4 Experiments
+III-E Comparison With Other Lightweight Forecasting Models
+Common lightweight models, such as DLinear [45] and LightTS [46], focus on the characteristics of the time series itself through different approaches.
+The main differences between TEFN and other lightweight models are as follows:
+  • Theoretical Foundation: While DLinear employs seasonal-trend decomposition and LightTS uses sampling techniques, evidence-based framework of TEFN fundamentally differs by modeling uncertainty through mass functions rather than direct feature extraction.
+This allows TEFN to handle ambiguous temporal patterns that challenge traditional decomposition methods.
+  • Feature Engineering: DLinear requires explicit decomposition of trends and seasonality, making it sensitive to decomposition accuracy.
+TEFN eliminates this need through its BPA module’s automatic uncertainty quantification, as shown in Equation (12).
+  • Computational Efficiency: Though all models emphasize efficiency, TEFN achieves parameter reduction through expectation fusion rather than architectural simplification.
+Expectation fusion directly adds the results of time BPA and channel BPA, without multiplication operation that the acceleration mode is similar to [50], and without introducing new modules.
+The parameter amount occupied in fusion operation is 0.
+Also, this will become a potential limitation, because in the engineering application scenarios of TEFN, due to the different characteristics of the data, other fusion methods may be used, such as concat and then use the linear link layer for fusion, or even use the attention mechanism with larger parameters for fusion.
+IV Experiments
 The experimental section aims to comprehensively evaluate TEFN’s performance in long-term time series forecasting.
@@
 Lastly, we will explore TEFN’s interpretability by analyzing the basic probability assignments generated by its BPA module.
-4.1 Dataset
+IV-A Dataset
 Our research employs the TEFN for the challenging task of long-term time series prediction.
@@
 These datasets encompass a wide range of real-world applications, providing a comprehensive evaluation of TEFN’s performance.
-The selected datasets are:
-  • Electricity [44]: This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time.
-It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
-  • ETT (4 subsets) [19]: This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges.
-It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
-  • Exchange [45]: This dataset captures exchange rate fluctuations between major currencies.
-It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
-  • Traffic [46]: This dataset contains traffic flow data collected from sensors in a city.
-It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
-  • Weather [47]: This dataset includes historical weather data, encompassing various meteorological parameters.
-It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
+The datasets are Electricity [51], ETT (4 subsets) [19], Exchange [52], Traffic [53], and Weather [54].
 Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities.
@@
 0
-4.2 Comparative Baseline
+IV-B Comparative Baseline
 To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models.
 These models represent various approaches and techniques commonly used in time series forecasting, providing a robust benchmark for assessing TEFN’s capabilities.
-The selected baseline models are as follows:
-  • Transformer-based Models:
-    • iTransformer (ICLR 2024) [48]: A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
-    • PatchTST (ICLR 2023) [36]: This model divides the time series into patches and applies transformer-based operations on each patch independently.
-It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
-    • Crossformer (ICLR 2023) [20]: This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
-    • FEDformer (ICML 2022) [52]: A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
-    • ETSformer [51]: A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
-    • Stationary (NeurIPS 2022) [42]: A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
-    • Autoformer (NeurIPS 2021) [45]: A self-supervised transformer model that learns to forecast future values without relying on explicit labels.
-It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
-  • Linear Models:
-    • Rlinear (ICLR 2024) [49]: This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data.
-It is effective at capturing periodic patterns and works well with longer input sequences.
-    • Dlinear (AAAI 2023) [39]: A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data.
-It serves as a simple yet effective baseline for comparison.
-    • LightTS [40]: A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
-  • Other Models:
-    • TiDE (ICLR 2023) [50]: A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
-    • TimesNet (ICLR 2023) [43]: A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
+The selected baseline models are as follows: Transformer-based Models, including iTransformer (ICLR 2024) [55], Crossformer (ICLR 2023) [20], FEDformer (ICML 2022) [59], ETSformer [58], Stationary (NeurIPS 2022) [48], and Autoformer (NeurIPS 2021) [52]; Linear Models, such as Rlinear (ICLR 2024) [56], DLinear (AAAI 2023) [45], and LightTS [46]; and Other Models, like TiDE (ICLR 2023) [57] and TimesNet (ICLR 2023).
 By comparing TEFN against this diverse set of baseline models, we can comprehensively assess its performance and identify its unique advantages and contributions to the field of time series forecasting.
-4.3 Implementation Details
+IV-C Implementation Details
 Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU.
@@
 For the sake of fairness, we used an open-source benchmark  [^1] The open-source benchmark Time-Series-Library located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations.
-The open-source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network..
+The repository address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network..
 The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam [61].
@@
 $MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (15)
-4.4 Forecasting Result
+IV-D Forecasting Result
 In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table II.
@@
 From the visualization results, TEFN fits the details and trends of time series better.
-4.5 Efficiency Comparison
+IV-E Efficiency Comparison
 TEFN is a compact model compared to the parameters of the Transformer.
@@
 It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices.
-As evidenced by the official code repository  [^2] https://github.com/ztxtech/Time-Evidence-Fusion-Network, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM.
+As evidenced by the code repository, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8 GB RAM.
 This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer [20] when deployed on portable computers.
@@
 The fitting of nonlinear functions also brings higher computational costs, therefore, in the case of efficiency priority, the linear version of TEFN is the preferred choice.
-4.6 Hyperparameter Sensitivity
-(a) $lr$ $|S|$MSE
-(b) Violin diagram of $|S|$ and MSE
-Figure 7: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
+IV-F Hyperparameter Sensitivity
 To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$.
@@
 It is worth noting that the order of magnitude of the MSE in Figure 8 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small.
-In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [55, 56]  [^3] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/).
-is conducted on prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 8.
-According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE.
-Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table III.
-The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation.
-Random parameter selection within the traversal range of TEFN yields stable prediction results.
-TABLE III: The variance of model error for traversal learning rate and sample space
+(a) lrlr - $|S|$ -MSE
+(b) Violin diagram of $|S|$ and MSE
+Figure 8: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
+TABLE III: Non-linearity of different datasets
+Non-linearity Dataset
+Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
+MSE 96 0.196 0.399 0.342 0.345 0.197 0.084 0.653 0.198
+192 0.195 0.442 0.452 0.382 0.275 0.172 0.602 0.239
+336 0.208 0.492 0.547 0.415 0.374 0.308 0.609 0.281
+720 0.244 0.516 0.763 0.484 0.451 0.792 0.649 0.346
+MAE 96 0.279 0.413 0.393 0.372 0.291 0.211 0.401 0.259
+192 0.282 0.435 0.455 0.391 0.352 0.308 0.375 0.302
+336 0.298 0.467 0.516 0.417 0.420 0.413 0.378 0.331
+720 0.331 0.511 0.630 0.465 0.458 0.669 0.399 0.384
+TABLE IV: The variance of model error for traversal learning rate and sample space
 Metric Dataset
-($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
+(VarVar) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
 MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
@@
 Figure 9: Heatmap of hyperparameter sensitivity correlation
-4.7 Ablation Study
-TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$.
-In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table IV.
+Moreover, to observe the relationship between the non-linearity and the performance of TEFN, we quantified the non-linearity of different datasets in an intuitive way: We directly use a linear layer to predict all the datasets.
+Since all the datasets have undergone the same standardization before prediction (following the settings of the original Benchmark), the errors are at the same level.
+Moreover, the larger the error, the greater the non-linearity.
+The nonlinear experimental results are in Table III.
+It can be found that the longer the prediction length, the greater the degree of nonlinearity.
+For different datasets, the error in the Traffic dataset is relatively large at different step lengths, showing a significant non-linearity.
+The non-linearity of the ETT dataset is relatively stable.
+In addition, Exchange dataset is greatly influenced by the prediction step length, reflecting that Exchange dataset shows linearity in the short term, but as the prediction step length increases, the linearity gradually decreases.
+In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [62, 63]  [^2] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/).
+is conducted on non-linearity, prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 9.
+According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE.
+In terms of nonlinearity, the forecasting error of the model and the non-linearity show a high positive correlation, such as the correlation between non-linearity MSE and MSE.
+This also means that TEFN will have a lower prediction error on datasets with low non-linearity, while the error will be relatively higher for datasets with high non-linearity.
+This comparison is only limited to the cross-dataset comparison of TEFN, and cannot reflect the performance comparison between different models.
+Building on this, a broader analysis offers insights into the performance of linear versus non-linear models.
+The observed positive correlation between non-linearity and prediction error appears to be a general phenomenon for linear models, as seen by comparing models like DLinear and RLinear in Table II with the non-linearity scores in Table III.
+While non-linear models can potentially mitigate this issue, their success is not guaranteed and often hinges on the alignment between their architectural design and the data’s specific characteristics.
+For instance, a model with a strong inductive bias for crossed variable relationships, such as Crossformer, may excel on one dataset such as Weather but not universally outperform simpler models on others.
+This highlights a crucial tradeoff.
+Non-linear models offer the potential for higher accuracy on complex data but at the cost of increased computational complexity, memory usage, and potentially lower robustness.
+Conversely, linear models like TEFN, while having a performance ceiling on highly non-linear data, provide significant benefits in terms of efficiency in Figure 7.
+In many scenarios, as evidenced by our results, they provide a highly competitive balance between performance and cost, making them a practical and effective choice for many real-world applications.
+Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table IV.
+The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation.
+On the other hand, we can observe that as the prediction step length increases, the variance also tends to increase, indicating that although non-linearity can improve the parameter sensitivity of the model, due to the fewer parameters of the linear model, its amplitude is limited.
+Random parameter selection within the traversal range of TEFN yields stable prediction results.
+IV-G Ablation Study
+TABLE V: The results of the ablation experiment
+Models TEFN w/o Norm w/o T w/o C Prob Concat Relu Tanh
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTm1 96 0.343 0.367 0.348 -1.44% 0.378 -2.91% 0.338 1.48% 0.365 0.55% 0.349 -1.72% 0.370 -0.81% 0.351 -2.28% 0.373 -1.61% 0.340 0.88% 0.367 0.00% 0.347 -1.15% 0.372 -1.34% 0.342 0.29% 0.369 -0.54%
+192 0.381 0.383 0.385 -1.04% 0.395 -3.04% 0.378 0.79% 0.383 0.00% 0.389 -2.06% 0.391 -2.05% 0.391 -2.56% 0.393 -2.54% 0.387 -1.55% 0.385 -0.52% 0.381 0.00% 0.385 -0.52% 0.381 0.00% 0.385 -0.52%
+336 0.414 0.404 0.428 -3.27% 0.428 -5.61% 0.410 0.98% 0.403 0.25% 0.422 -1.90% 0.413 -2.18% 0.422 -1.90% 0.413 -2.18% 0.414 0.00% 0.402 0.50% 0.415 -0.24% 0.406 -0.49% 0.409 1.22% 0.404 0.00%
+720 0.475 0.438 0.471 0.85% 0.447 -2.01% 0.474 0.21% 0.438 0.00% 0.484 -1.86% 0.447 -2.01% 0.489 -2.86% 0.449 -2.45% 0.476 -0.21% 0.438 0.00% 0.477 -0.42% 0.440 -0.45% 0.473 0.42% 0.438 0.00%
+ETTm2 96 0.181 0.264 0.185 -2.16% 0.277 -4.69% 0.181 0.00% 0.264 0.00% 0.181 0.00% 0.263 0.38% 0.183 -1.09% 0.265 -0.38% 0.198 -8.59% 0.281 -6.05% 0.182 -0.55% 0.264 0.00% 0.181 0.00% 0.263 0.38%
+192 0.246 0.304 0.284 -13.38% 0.352 -13.64% 0.247 -0.40% 0.304 0.00% 0.246 0.00% 0.304 0.00% 0.246 0.00% 0.304 0.00% 0.247 -0.40% 0.305 -0.33% 0.247 -0.40% 0.304 0.00% 0.246 0.00% 0.304 0.00%
+336 0.307 0.343 0.372 -17.47% 0.416 -17.55% 0.307 0.00% 0.343 0.00% 0.307 0.00% 0.342 0.29% 0.307 0.00% 0.343 0.00% 0.308 -0.32% 0.343 0.00% 0.308 -0.32% 0.343 0.00% 0.308 -0.32% 0.344 -0.29%
+720 0.407 0.398 0.597 -31.83% 0.533 -25.33% 0.406 0.25% 0.399 -0.25% 0.407 0.00% 0.398 0.00% 0.407 0.00% 0.398 0.00% 0.405 0.49% 0.398 0.00% 0.407 0.00% 0.399 -0.25% 0.407 0.00% 0.400 -0.50%
+ETTh1 96 0.383 0.391 0.382 0.26% 0.394 -0.76% 0.383 0.00% 0.392 -0.26% 0.384 -0.26% 0.392 -0.26% 0.387 -1.03% 0.394 -0.76% 0.385 -0.52% 0.393 -0.51% 0.383 0.00% 0.393 -0.51% 0.393 -2.54% 0.400 -2.25%
+192 0.433 0.419 0.474 -8.65% 0.472 -11.23% 0.432 0.23% 0.420 -0.24% 0.435 -0.46% 0.422 -0.71% 0.436 -0.69% 0.422 -0.71% 0.431 0.46% 0.419 0.00% 0.432 0.23% 0.422 -0.71% 0.432 0.23% 0.421 -0.48%
+336 0.475 0.441 0.477 -0.42% 0.452 -2.43% 0.476 -0.21% 0.440 0.23% 0.478 -0.63% 0.444 -0.68% 0.481 -1.25% 0.445 -0.90% 0.475 0.00% 0.440 0.23% 0.472 0.64% 0.441 0.00% 0.473 0.42% 0.440 0.23%
+720 0.475 0.464 0.505 -5.94% 0.499 -7.01% 0.474 0.21% 0.461 0.65% 0.478 -0.63% 0.468 -0.85% 0.481 -1.25% 0.469 -1.07% 0.479 -0.84% 0.462 0.43% 0.472 0.64% 0.464 0.00% 0.470 1.06% 0.462 0.43%
+ETTh2 96 0.288 0.337 0.345 -16.52% 0.389 -13.37% 0.287 0.35% 0.337 0.00% 0.289 -0.35% 0.337 0.00% 0.289 -0.35% 0.337 0.00% 0.290 -0.69% 0.339 -0.59% 0.306 -5.88% 0.351 -3.99% 0.293 -1.71% 0.342 -1.46%
+192 0.375 0.392 0.557 -32.68% 0.499 -21.44% 0.376 -0.27% 0.393 -0.25% 0.376 -0.27% 0.390 0.51% 0.377 -0.53% 0.390 0.51% 0.377 -0.53% 0.395 -0.76% 0.376 -0.27% 0.394 -0.51% 0.375 0.00% 0.392 0.00%
+336 0.423 0.434 0.588 -28.06% 0.534 -18.73% 0.416 1.68% 0.430 0.93% 0.409 3.42% 0.423 2.60% 0.415 1.93% 0.426 1.88% 0.416 1.68% 0.430 0.93% 0.421 0.48% 0.433 0.23% 0.420 0.71% 0.432 0.46%
+720 0.434 0.446 0.895 -51.51% 0.662 -32.63% 0.458 -5.24% 0.463 -3.67% 0.429 1.17% 0.445 0.22% 0.420 3.33% 0.439 1.59% 0.498 -12.85% 0.485 -8.04% 0.470 -7.66% 0.469 -4.90% 0.450 -3.56% 0.457 -2.41%
+Electricity 96 0.197 0.273 0.194 1.55% 0.277 -1.44% 0.197 0.00% 0.276 -1.09% 0.197 0.00% 0.274 -0.36% 0.198 -0.51% 0.274 -0.36% 0.197 0.00% 0.274 -0.36% 0.197 0.00% 0.274 -0.36% 0.197 0.00% 0.273 0.00%
+192 0.197 0.276 0.193 2.07% 0.280 -1.43% 0.195 1.03% 0.275 0.36% 0.197 0.00% 0.277 -0.36% 0.197 0.00% 0.276 0.00% 0.197 0.00% 0.276 0.00% 0.197 0.00% 0.277 -0.36% 0.196 0.51% 0.277 -0.36%
+336 0.212 0.292 0.206 2.91% 0.295 -1.02% 0.209 1.44% 0.290 0.69% 0.211 0.47% 0.291 0.34% 0.212 0.00% 0.292 0.00% 0.211 0.47% 0.292 0.00% 0.211 0.47% 0.291 0.34% 0.211 0.47% 0.292 0.00%
+720 0.253 0.325 0.242 4.55% 0.329 -1.22% 0.252 0.40% 0.324 0.31% 0.253 0.00% 0.324 0.31% 0.253 0.00% 0.324 0.31% 0.253 0.00% 0.325 0.00% 0.254 -0.39% 0.326 -0.31% 0.252 0.40% 0.324 0.31%
+Exchange 96 0.082 0.198 0.078 5.13% 0.201 -1.49% 0.086 -4.65% 0.203 -2.46% 0.081 1.23% 0.197 0.51% 0.083 -1.20% 0.198 0.00% 0.082 0.00% 0.198 0.00% 0.085 -3.53% 0.203 -2.46% 0.088 -6.82% 0.206 -3.88%
+192 0.176 0.297 0.169 4.14% 0.304 -2.30% 0.202 -12.87% 0.320 -7.19% 0.172 2.33% 0.293 1.37% 0.176 0.00% 0.296 0.34% 0.177 -0.56% 0.299 -0.67% 0.180 -2.22% 0.301 -1.33% 0.174 1.15% 0.297 0.00%
+336 0.328 0.412 0.269 21.93% 0.396 4.04% 0.336 -2.38% 0.417 -1.20% 0.323 1.55% 0.409 0.73% 0.320 2.50% 0.408 0.98% 0.324 1.23% 0.409 0.73% 0.340 -3.53% 0.420 -1.90% 0.330 -0.61% 0.412 0.00%
+720 0.861 0.700 0.735 17.14% 0.674 3.86% 0.894 -3.69% 0.713 -1.82% 0.852 1.06% 0.694 0.86% 0.971 -11.33% 0.736 -4.89% 0.918 -6.21% 0.721 -2.91% 0.912 -5.59% 0.722 -3.05% 0.899 -4.23% 0.715 -2.10%
+Traffic 96 0.645 0.383 0.652 -1.07% 0.397 -3.53% 0.646 -0.15% 0.385 -0.52% 0.647 -0.31% 0.385 -0.52% 0.645 0.00% 0.384 -0.26% 0.646 -0.15% 0.388 -1.29% 0.647 -0.31% 0.386 -0.78% 0.644 0.16% 0.379 1.06%
+192 0.598 0.360 0.598 0.00% 0.370 -2.70% 0.597 0.17% 0.360 0.00% 0.597 0.17% 0.359 0.28% 0.598 0.00% 0.361 -0.28% 0.598 0.00% 0.358 0.56% 0.598 0.00% 0.360 0.00% 0.596 0.34% 0.356 1.12%
+336 0.605 0.362 0.605 0.00% 0.373 -2.95% 0.604 0.17% 0.363 -0.28% 0.605 0.00% 0.363 -0.28% 0.606 -0.17% 0.362 0.00% 0.605 0.00% 0.362 0.00% 0.605 0.00% 0.364 -0.55% 0.603 0.33% 0.359 0.84%
+720 0.644 0.382 0.645 -0.16% 0.395 -3.29% 0.642 0.31% 0.382 0.00% 0.643 0.16% 0.381 0.26% 0.643 0.16% 0.381 0.26% 0.644 0.00% 0.383 -0.26% 0.648 -0.62% 0.389 -1.80% 0.641 0.47% 0.382 0.00%
+Weather 96 0.182 0.227 0.180 1.11% 0.245 -7.35% 0.187 -2.67% 0.230 -1.30% 0.196 -7.14% 0.235 -3.40% 0.194 -6.19% 0.234 -2.99% 0.183 -0.55% 0.228 -0.44% 0.184 -1.09% 0.230 -1.30% 0.183 -0.55% 0.228 -0.44%
+192 0.227 0.262 0.221 2.71% 0.288 -9.03% 0.227 0.00% 0.261 0.38% 0.241 -5.81% 0.271 -3.32% 0.240 -5.42% 0.270 -2.96% 0.227 0.00% 0.262 0.00% 0.230 -1.30% 0.265 -1.13% 0.227 0.00% 0.263 -0.38%
+336 0.279 0.298 0.275 1.45% 0.330 -9.70% 0.281 -0.71% 0.300 -0.67% 0.292 -4.45% 0.306 -2.61% 0.290 -3.79% 0.305 -2.30% 0.282 -1.06% 0.302 -1.32% 0.282 -1.06% 0.301 -1.00% 0.278 0.36% 0.298 0.00%
+720 0.352 0.344 0.334 5.39% 0.375 -8.27% 0.351 0.28% 0.344 0.00% 0.364 -3.30% 0.354 -2.82% 0.365 -3.56% 0.353 -2.55% 0.352 0.00% 0.345 -0.29% 0.351 0.28% 0.344 0.00% 0.351 0.28% 0.344 0.00%
+We conducted experiments on the components of TEFN, the replacement of the probability layer, the concat fusion method, and different BPA nonlinear implementations in Table V.
+At the same time, we calculated the error change ratio $\gamma$ compared with the original TEFN in Equation (16).
+$\gamma=\frac{\text{Metric}_{TEFN}-\text{Metric}}{\text{Metric}}*100\%$ (16)
+IV-G1 Components
+TEFN consists of a norm layer, a time dimension projection layer Norm, a time fusion layer T, and a channel fusion layer C.
 Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error.
@@
 It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
-We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting.
-We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table IV.
-In the aspect of all the results of ablation, TEFN can effectively improve performance by combining the dimensions of time $T$ and channel $C$ in most cases.
 There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored.
-According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
+The forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
 $l_{C+T}\leqslant min(l_{C},l_{T})$ (17)
-TABLE IV: The results of the ablation experiment
-Metric Target Model Dataset
-Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
-MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
-C 0.196$\downarrow$ 0.386$\uparrow$ 0.288 0.339$\downarrow$ 0.185$\uparrow$ 0.083$\uparrow$ 0.644$\downarrow$ 0.185$\uparrow$
-T 0.197 0.386$\uparrow$ 0.291$\uparrow$ 0.356$\uparrow$ 0.182$\uparrow$ 0.083$\uparrow$ 0.645 0.195$\uparrow$
-192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
-C 0.196$\downarrow$ 0.435$\uparrow$ 0.375 0.382$\uparrow$ 0.246 0.179$\uparrow$ 0.597$\downarrow$ 0.230$\uparrow$
-T 0.197 0.435$\uparrow$ 0.375 0.390$\uparrow$ 0.247$\uparrow$ 0.175$\downarrow$ 0.598 0.240$\uparrow$
-336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
-C 0.210$\downarrow$ 0.479$\uparrow$ 0.423 0.411$\downarrow$ 0.307 0.336$\uparrow$ 0.604$\downarrow$ 0.280$\uparrow$
-T 0.212 0.479$\uparrow$ 0.414$\downarrow$ 0.423$\uparrow$ 0.307 0.324$\downarrow$ 0.606$\uparrow$ 0.291$\uparrow$
-720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
-C 0.253 0.478$\uparrow$ 0.457$\uparrow$ 0.472$\downarrow$ 0.407 0.947$\uparrow$ 0.642$\downarrow$ 0.354$\uparrow$
-T 0.254$\uparrow$ 0.479$\uparrow$ 0.442$\uparrow$ 0.487$\uparrow$ 0.408$\uparrow$ 0.871$\uparrow$ 0.643$\downarrow$ 0.364$\uparrow$
-Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
-C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
-T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
-MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
-C 0.273$\downarrow$ 0.394$\uparrow$ 0.339$\uparrow$ 0.366$\downarrow$ 0.269$\uparrow$ 0.199$\uparrow$ 0.384 0.228$\uparrow$
-T 0.274 0.393$\uparrow$ 0.339$\uparrow$ 0.377$\uparrow$ 0.265$\uparrow$ 0.199$\uparrow$ 0.383$\downarrow$ 0.234$\uparrow$
-192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
-C 0.275$\downarrow$ 0.421$\uparrow$ 0.393$\uparrow$ 0.385$\uparrow$ 0.305$\uparrow$ 0.300$\uparrow$ 0.361$\uparrow$ 0.264$\uparrow$
-T 0.276 0.421$\uparrow$ 0.390$\downarrow$ 0.392$\uparrow$ 0.305$\uparrow$ 0.296$\downarrow$ 0.361$\uparrow$ 0.270$\uparrow$
-336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
-C 0.291$\downarrow$ 0.443$\uparrow$ 0.434$\downarrow$ 0.403$\downarrow$ 0.343 0.419$\uparrow$ 0.363 0.298
-T 0.292 0.445$\uparrow$ 0.425$\downarrow$ 0.413$\uparrow$ 0.342$\downarrow$ 0.410$\downarrow$ 0.362$\downarrow$ 0.306$\uparrow$
-720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
-C 0.325 0.463$\downarrow$ 0.462$\uparrow$ 0.436$\downarrow$ 0.399 0.739$\uparrow$ 0.382 0.347$\uparrow$
-T 0.324$\downarrow$ 0.469$\uparrow$ 0.454$\uparrow$ 0.449$\uparrow$ 0.398$\downarrow$ 0.702$\uparrow$ 0.382 0.354$\uparrow$
-Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
-C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
-T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
-$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
-(a) Channel BPA
-(b) Time BPA
-Figure 9: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time.
-The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
-4.8 Robustness Experiments
+Furthermore, we empirically demonstrated that removing the normalization layer led to a substantial performance improvement on certain datasets, such as the Exchange dataset.
+The normalization layer typically accelerates the model’s training process; however, it also presents issues.
+During denormalization, it relies solely on the standard deviation and mean of the input samples, overlooking potential changes in these statistics.
+Conversely, the normalization layer proves beneficial for the ETT dataset.
+Without it, the model exhibits severe underfitting, as evidenced by a twofold increase in the MSE for ETTh2-720.
+IV-G2 Probability Layer
+To elucidate the disparity between the Basic Probability Assignment (BPA) and a traditional probability model, we conducted an experiment by removing two BPA layers.
+As depicted in Equation (12), when the mass function degenerates into a probability, the BPA module simplifies to an ordinary linear layer.
+At this juncture, the time dimension projection accomplishes an expectation prediction from a probabilistic perspective.
+Moreover, BPA can be regarded as an imprecise estimate of probability, as it takes into account a broader spectrum of possibilities.
+Consequently, the output of the BPA module increases the dimensionality of the data.
+Empirical results indicate that, in comparison to the original TEFN, most of the forecasting errors have risen.
+IV-G3 Concat Fusion Method
+The integration approach adopted by TEFN is expectation fusion.
+Therefore, we can directly add the expanded dimensions of the two BPAs to complete the prediction without introducing additional fusion layers.
+Here, we use the concat method for comparison.
+We concatenate the results of the time BPA and the channel BPA, and then use a linear layer for the final output.
+Experimental data shows that the concat method does not significantly improve the prediction performance.
+However, due to the concat operation and the addition of the fusion layer, new parameters are introduced, and more memory is occupied.
+IV-G4 Nonlinear Implementation of BPA
+We added a nonlinear activation layer after the BPA module to enhance the nonlinear fitting ability of TEFN using the residual method described in Equation (18).
+The ReLU function, commonly used in classification tasks, and the Tanh activation function, commonly used in time series analysis, have been selected here.
+Experimental data indicates that the performance of ReLU is not ideal because ReLU essentially filters out weights less than 0, rather than performing a continuous transformation.
+In contrast, Tanh improves the forecasting performance of TEFN in most scenarios because it is a smooth function suitable for time series prediction.
+$x\leftarrow x+\text{act}(x)$ (18)
+IV-H Robustness Experiments
 TEFN is a highly robust neural network.
 To verify the robustness of TEFN in real prediction, a comparative experiment was conducted by adding noise to the dataset.
-The unit noise $\gamma$ follows the Gaussian distribution $N(0,1)$ with an expectation of 00 and a variance of $1$.
+The unit noise $\varepsilon$ follows the Gaussian distribution $N(0,1)$ with an expectation of 0 and a variance of $1$.
 Assume that the input time series is $x$, with time and channel dimensions.
-To adapt to the scale of the data, the transformation as Eq.(18) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function.
+To adapt to the scale of the data, the transformation as Equation (19) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function.
 So, the noisy data experiments are compared with the original data experiments.
-$x^{\prime}_{ij}=x_{ij}+std(x_{i})*\gamma$ (18)
-The experimental results of robustness testing are shown in the Table V.
+$x^{\prime}_{ij}=x_{ij}+\text{std}(x_{i})*\varepsilon$ (19)
+The experimental results of robustness testing are shown in the Table VI.
 Continuing from the analysis, the experimental outcomes underscore the robust nature of TEFN, where the introduction of Gaussian noise to the dataset led to only marginal increases in prediction errors, with most variations being insignificant relative to the original errors.
 Notably, in long-term forecasting scenarios, a counterintuitive improvement in accuracy was observed in some instances, indicating that TEFN’s adaptability and learning capacity in complex, dynamic environments extend beyond mere resilience to noise, potentially leveraging subtle perturbations to enhance prediction performance.
-TABLE V: The results of robustness experiments
+TABLE VI: The results of robustness experiments
 Dataset Pred Original Noisy
@@
 336 0.414 0.404 0.499 ($\uparrow$ 0.085) 0.458 ($\uparrow$ 0.054)
-720 0.475 0.438 0.381 ($\downarrow$ 0.094) 0.398 ($\downarrow$ 0.040)
+720 0.475 0.438 0.381 ($\downarrow$ 0.094) 0.398 ($\downarrow$ 0.04)
 ETTm2 96 0.181 0.264 0.254 ($\uparrow$ 0.073) 0.312 ($\uparrow$ 0.048)
 192 0.246 0.304 0.314 ($\uparrow$ 0.068) 0.350 ($\uparrow$ 0.046)
-336 0.307 0.343 0.411 ($\uparrow$ 0.104) 0.403 ($\uparrow$ 0.060)
+336 0.307 0.343 0.411 ($\uparrow$ 0.104) 0.403 ($\uparrow$ 0.06)
 720 0.407 0.398 0.199 ($\downarrow$ 0.208) 0.281 ($\downarrow$ 0.117)
@@
 720 0.644 0.382 0.774 ($\uparrow$ 0.130) 0.494 ($\uparrow$ 0.112)
-Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.050) 0.269 ($\uparrow$ 0.042)
+Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.05) 0.269 ($\uparrow$ 0.042)
 192 0.227 0.262 0.285 ($\uparrow$ 0.058) 0.305 ($\uparrow$ 0.043)
-336 0.279 0.298 0.357 ($\uparrow$ 0.078) 0.350 ($\uparrow$ 0.052)
+336 0.279 0.298 0.357 ($\uparrow$ 0.078) 0.35 ($\uparrow$ 0.052)
 720 0.352 0.344 0.190 ($\downarrow$ 0.162) 0.237 ($\downarrow$ 0.107)
@@
 720 0.861 0.700 0.092 ($\downarrow$ 0.769) 0.213 ($\downarrow$ 0.487)
-4.9 Interpretability Analysis
+IV-I Interpretability Analysis
+(a) Channel BPA
+(b) Time BPA
+Figure 10: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time.
+The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
 The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable.
@@
 By analyzing the shape and parameters of these membership functions, users can gain a clear understanding of the model’s perception of data features and their influence on predictions.
-For instance, it becomes evident which features are deemed more important than others, and how the model responds to variations within these features.
+For instance, it becomes clear which features are deemed more important than others, and how the model responds to variations within these features.
 This visualization not only makes the model’s inner workings more transparent but also allows users to intuitively assess the model’s confidence in its predictions and the level of uncertainty associated with each outcome.
@@
 By exploring these aspects, users can develop a comprehensive understanding of the model’s decision-making process and prediction outcomes, fostering trust and confidence in the model’s capabilities.
-4.10 Limitation and Future Prospects
+IV-J Limitation
 TEFN is a naive and unmodified model.
@@
 Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well.
+The experiment shows that adding a nonlinear activation function can effectively improve the experimental performance of TEFN.
 In addition, for the fusion of mass distribution in the face of the failure of the classic DSR for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern.
@@
 Effective parameter initialization may exist, for example, initializing corresponding parameters in a non-neural network way may lead to faster convergence.
-5 Conclusion
+V Conclusion
 This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets.